### v1

# Quantum and Classical Communication Complexity of Permutation-Invariant Functions

Abstract
This paper gives a nearly tight characterization of the quantum communication complexity of the permutation-invariant Boolean functions. With such a characterization, we show that the quantum and randomized communication complexity of the permutation-invariant Boolean functions are quadratically equivalent (up to a logarithmic factor). Our results extend a recent line of research regarding query complexity [2, 17, 11] to communication complexity, showing symmetry prevents exponential quantum speedups.
Furthermore, we show the Log-rank Conjecture holds for any non-trivial total permutation-invariant Boolean function. Moreover, we establish a relationship between the quantum/classical communication complexity and the approximate rank of permutation-invariant Boolean functions. This implies the correctness of the Log-approximate-rank Conjecture for permutation-invariant Boolean functions in both randomized and quantum settings (up to a logarithmic factor).
1 Introduction
Exploring quantum advantages is a key problem in the realm of quantum computing. Numerous work focuses on analyzing and characterizing quantum advantages, such as [6, 14, 25, 21, 29, 47]. It has been known that quantum computing demonstrates a potential exponential speedup to solve certain problems than classical computers, such as Simon’s problem [43] and integer factoring problem [42]. However, for some problems, the quantum speedups can be at most polynomial, including the unstructured search problems [26] and collision finding problems [4]. In light of the aforementioned phenomenon, Aaronson [1] proposed such a problem: How much structure is needed for huge quantum speedups?
Regarding the above problem, there exist two major directions to explore the structure needed for quantum speedups in the query model, which is a complexity model commonly used to describe quantum advantages. On the one hand, Aaronson and Ambainis [2] conjectured the acceptance probability of a quantum query algorithm to compute a Boolean function can be approximated by a classical deterministic algorithm with only a polynomial increase in the number of queries, which is still one of most important conjecture in the field of Boolean analysis. On the other hand, Watrous conjectured that the quantum and randomized query complexities are also polynomially equivalent for any permutation-invariant function [2]. Along this direction, Aaronson and Ambainis [2] initiated the study on the quantum speedup of permutation-invariant functions with respect to query complexity. They demonstrated that a function invariant under full symmetry does not exhibit exponential quantum speedups, even if the function is partial, thereby resolving the Watrous conjecture. (Interested readers may refer to [2] for a more detailed introduction.) Furthermore, Chailloux [17] expanded upon their work by providing a tighter bound and removing a technical dependence of output symmetry. Recently, Ben-David, Childs, Gilyén, Kretschmer, Podder and Wang [11] further proved that hypergraph symmetries in the adjacency matrix model allow at most a polynomial separation between randomized and quantum query complexities. All the above results demonstrated that symmetries break exponential quantum speedups in the query model.
While the study of problem structure in the roles of quantum speedups has obtained considerable attention in the query model, it is a natural question to consider whether we can derive similar results in other computation models. The communication complexity model comes to attention as it is also extensively used to demonstrate quantum advantages. Furthermore, while the exponential gap between quantum and classical communication models has been shown in many works [38, 7, 22, 23, 36], there are also some problems in communication models that demonstrate at most polynomial quantum speedups, such as set disjointness problem [39] and (gap) Hamming distance problem [28, 41, 46, 19]. Therefore, it is a meaningful question to consider how much structure is needed for significant quantum speedups in the communication complexity model. More specifically, while symmetry breaks quantum exponential advantages in the query model, does there exist a similar conclusion in the communication complexity model? In this paper, we investigate a variant of the Watrous conjecture concerning the quantum and randomized communication complexities of permutation-invariant functions as follows. Briefly, a permutation-invariant Boolean function is a function that is invariant under permutations of its inputs (see Definition 2.1 for a formal definition).
Conjecture 1.1 (Communication complexity version of the Watrous Conjecture.).
For any permutation-invariant function $f:[m]^{n}\times[m]^{n}\rightarrow\left\{-1,1,*\right\}$, $R(f)\leq Q^{*}(f)^{O(1)}$, where $R(f)$ and $Q^{*}(f)$ are the randomized and quantum communication complexities of $f$, respectively.
Furthermore, we study the Log-rank Conjecture proposed by Lovasz and Saks [34], a long-standing open problem in communication complexity. Despite its slow progress on total functions [12, 35, 31], the conjecture has been shown for several subclasses of total permutation-invariant Boolean functions [16] and $\mathsf{XOR}$-symmetric functions [48]. Lee and Shraibman further proposed the Log-approximate-rank Conjecture, stating that the randomized communication and the logarithm of the approximate rank of the input matrix are polynomially equivalent. Surprisingly, this conjecture was later proven false [20], even for its quantum counterpart [5, 44].
In this paper, we investigate both conjectures for permutation-invariant functions.
Conjecture 1.2 (Log-rank Conjecture for permutation-invariant functions.).
For any total permutation-invariant function $f:\{0,1\}^{n}\times\{0,1\}^{n}\rightarrow\left\{-1,1\right\}$, $D(f)\leq\left(\log\operatorname{\mathrm{rank}}(f)\right)^{O(1)}$, where $\operatorname{\mathrm{rank}}(f)$ is the rank of the input matrix of $f$.
Conjecture 1.3 (Log-Approximate-Rank Conjecture for permutation-invariant functions.).
For any (total or partial) permutation-invariant function $f:\{0,1\}^{n}\times\{0,1\}^{n}\rightarrow\left\{-1,1,*\right\}$, $R(f)\leq\left(\log\operatorname{\widetilde{rank}}(f)\right)^{O(1)}$, where $\operatorname{\widetilde{rank}}(f)$ is the approximate rank of the input matrix of $f$ (see Definition 2.4 for a formal definition).
Conjecture 1.4 (Quantum Log-Approximate-Rank Conjecture for permutation-invariant functions.).
For any (total or partial) permutation-invariant function $f:\{0,1\}^{n}\times\{0,1\}^{n}\rightarrow\left\{-1,1,*\right\}$, $Q(f)\leq\left(\log\operatorname{\widetilde{rank}}(f)\right)^{O(1)}$.
1.1 Our Contribution
To study the communication complexity version of the Watrous conjecture, we start with permutation-invariant Boolean functions, which are essential to analyze general permutation-invariant functions. We show that for any permutation-invariant Boolean function, its classical communication complexity has at most a quasi-quadratic blowup comparing to its quantum communication complexity (Theorem 1.5). Thus, we cannot hope for exponential quantum speedups of permutation-invariant Boolean functions. Additionally, Theorem 1.5 gives a nearly tight bound on the quantum communication complexity. Furthermore, we show that every non-trivial permutation-invariant Boolean function satisfies the Log-rank Conjecture in Theorem 1.6. To resolve the (quantum) Log-Approximate-Rank Conjecture, we investigate the relationship between the quantum/classical communication complexities and the approximate rank of any permutation-invariant Boolean function in Theorem 1.7.
Consider a Boolean function $f$. Let $D(f),R(f),Q(f)$ and $Q^{*}(f)$ be the deterministic communication complexity, randomized communication complexity, quantum communication complexity without prior entanglement, and quantum communication complexity of $f$, respectively. Let $\operatorname{\mathrm{rank}}\left(f\right)$ and $\operatorname{\widetilde{rank}}\left(f\right)$ be the rank and approximate rank of $f$. We summarize our results below [^1] In Theorems 1.5 and 1.7, $\widetilde{O}\left(M\left(f\right)\right)=O\left(M(f)\log^{2}n\log\log n\right)$ for any complexity measure $M$..
Theorem 1.5.
For any permutation-invariant function $f:\{0,1\}^{n}\times\{0,1\}^{n}\to\{-1,1,*\}$ in Definition 2.1, the followings hold:
$\Omega\left(m\left(f\right)\right)\leq R(f)\leq\widetilde{O}\left(m\left(f\right)^{2}\right)\leq\widetilde{O}\left(Q^{*}(f)^{2}\right)\text{ and }$
$\Omega\left(m\left(f\right)\right)\leq Q^{*}(f)\leq Q(f)\leq\widetilde{O}\left(m\left(f\right)\right),$
where $m(f)$ is a measure defined in Definition 3.1. Hence, $R(f)\leq\widetilde{O}(Q^{*}(f)^{2})$ for any permutation-invariant function $f$.
The complexity measure $m(\cdot)$ is inspired by the work [24], where Gahzi et al. introduced a complexity measure to capture $R(f)$. It is worth noting that their complexity measure is equivalent up to a fourth power of $R(f)$, while our complexity measure $m(\cdot)$ is quadratically related to $R(f)$ and almost tightly characterizes the quantum communication complexity.
Theorem 1.6.
For any non-trivial total permutation-invariant function $f:\{0,1\}^{n}\times\{0,1\}^{n}\to\{-1,1\}$ in Definition 2.1, we have
$D(f)=O\left(\log^{2}\operatorname{\mathrm{rank}}\left(f\right)\right).$
Theorem 1.7.
For any permutation-invariant function $f:\{0,1\}^{n}\times\{0,1\}^{n}\to\{-1,1,*\}$ in Definition 2.1, we have
$R(f)=\widetilde{O}\left(\log^{2}\operatorname{\widetilde{rank}}\left(f\right)\right)~{}\text{ and}~{}Q(f)=\widetilde{O}\left(\log\operatorname{\widetilde{rank}}\left(f\right)\right).$
1.2 Proof Techniques
In this section, we give a high-level technical overview of our main results.
1.2.1 Lower Bound
We outline our approaches to obtain the lower bound on the quantum communication complexity, rank and approximate rank of permutation-invariant functions below:
  1. 1.
    Quantum communication complexity and approximate rank: In Theorem 1.5, to prove $Q^{*}(f)=\Omega(m(f))$ for any permutation-invariant function $f$, we use the following two-step reduction (see Lemmas 3.4 and 3.2): First, we reduce the lower bound of the quantum communication complexity of the Exact Set-Inclusion Problem ($\mathsf{ESetInc}$, Definition 2.3) to Paturis’s approximate degree of symmetric functions [37] by the pattern matrix method [40], a well-known method for lower bound analysis in quantum communication complexity. Second, we reduce the lower bound of any permutation-invariant function to the lower bound of $\mathsf{ESetInc}$. In Theorem 1.7, we use a similar method to prove the lower bound of approximate rank: $\log\operatorname{\widetilde{rank}}\left(f\right)=\Omega(m(f))$.
  2. 2.
    Rank: In Theorem 1.6, we reduce the lower bound of the rank of total permutation-invariant functions to the lower bound of the rank of some representative function instances, such as the set disjointness problem and the equality problem (see Lemma 4.1).
1.2.2 Upper Bound
We use the following methods to show the upper bounds on the communication complexity of permutation-invariant functions in the randomized, quantum, and deterministic models.
  1. 1.
    Randomized and quantum models: In Theorem 1.5, to prove $R(f)\leq\widetilde{O}\left(m\left(f\right)^{2}\right)$ for any permutation-invariant function $f$, we first propose a randomized protocol to solve the Set-Inclusion problem ($\mathsf{SetInc}$, Definition 2.3) using a well-suited sampling method according to the parameters of $\mathsf{SetInc}$ (see Lemma 3.9). Afterward, we use this protocol as a subroutine to solve any permutation-invariant function based on binary search (see Theorem 3.3). Furthermore, to prove $Q(f)\leq\widetilde{O}\left(m\left(f\right)\right)$, we use the quantum amplitude amplification technique [13, 27] to speed up the above randomized protocol to solve $\mathsf{SetInc}$ (see Lemma 3.10).
  2. 2.
    Deterministic model: In Theorem 1.6, to give an upper bound of deterministic communication complexity of total permutation-invariant functions, we propose a deterministic protocol as follows (see Lemma 4.2): Alice and Bob first share their Hamming weight of inputs, and decide who sends the input to the other party according to the definition of function and the Hamming weight of inputs. The party that has all the information of inputs will output the answer. Combining Lemmas 4.1 (described in Section 1.2.1) and 4.2, Theorem 1.6 can be proved.
1.3 Related Work
The need for structure in quantum speedups has been studied in the query model extensively. Beals, Buhrman, Cleve, Mosca and de Wolf [8] demonstrated that there exists at most a polynomial quantum speedup for total Boolean functions in the query model. Moreover, Aaronson and Ambainis [2] established that even partial symmetric functions do not allow super-polynomial quantum speedups. Chailloux [17] further improved this result to a broader class of symmetric functions. Ben-David, Childs, Gilyén, Kretschmer, Podder and Wang [11] later analyzed the quantum advantage for functions that are symmetric under different group actions systematically. Ben-David [10] established a quantum and classical polynomial equivalence for a certain set of functions satisfying a specific symmetric promise. Aaronson and Ben-David [3] proved that if domain $D$ satisfies $D=\text{poly}(n)$, there are at most polynomial quantum speedups for computing an $n$-bit partial Boolean function.
In terms of communication complexity, there are a few results that imply the polynomial equivalence between quantum and classical communication complexity for several instances of permutation-invariant functions. Examples include $\mathsf{AND}$-symmetric functions [39], Hamming distance problem [28, 18], $\mathsf{XOR}$-symmetric functions [48]. While the above results characterized quantum advantage for a certain class of permutation-invariant Boolean functions, our work provides a systemic analysis of all permutation-invariant Boolean functions.
The study of the Log-rank Conjecture and the Log-Approximate-Rank Conjecture has a rich history. Here, we only survey the results about the Log-rank Conjecture and the Log-Approximate-Rank Conjecture about permutation-invariant Boolean functions. Buhuman and de Wolf [16] verified the correctness of the Log-rank Conjecture for $\mathsf{AND}$-symmetric functions. Combining the results of Razborov [39], Sherstov [40] and Suruga [45], it is implied that the Log-Approximate-Rank Conjecture holds for $\mathsf{AND}$-symmetric functions both in the randomized and quantum settings. Moreover, the result of Zhang and Shi [48] implies the Log-rank Conjecture and the (quantum) Log-Approximate-Rank Conjecture hold for $\mathsf{XOR}$-symmetric functions.
In previous work, Ghazi, Kamath and Sudan [24] introduced a complexity measure, which is polynomially equivalent to the randomized communication complexity of permutation-invariant functions defined in Definition 2.1. This paper is inspired by their work.
1.4 Organization
The remaining part of the paper is organized as follows. In Section 2, we state some notations and definitions used in this paper. In Section 3, we study the quantum and classical communication complexities of permutation-invariant functions. In Section 4, we show the Log-rank Conjecture holds for non-trivial total permutation-invariant functions. In Section 5, we study the Log-approximate Conjecture of permutation-invariant functions both in quantum and classical setting. Finally, a conclusion is made in Section 6. The appendices contain a section on extended preliminaries and omitted proofs.
2 Preliminaries
We introduce the notations and definitions used in this paper.
[…]
2.1 Boolean Functions
A partial function is a function defined only on a subset of its domain $\mathcal{X}$. Formally, given a partial Boolean function $f\colon\mathcal{X}\rightarrow\left\{-1,1,*\right\}$, $f(x)$ is undefined for $x\in\mathcal{X}$ if $f(x)=*$. A total function is a function that is defined on the entire domain. We say $f\colon\mathcal{X}\rightarrow\left\{-1,1,*\right\}$ is a subfunction of $g\colon\mathcal{X}\rightarrow\left\{-1,1,*\right\}$ if $f(x)=g(x)$ or $f(x)=*$ for any $x\in\mathcal{X}$.
A Boolean predicate is a partial function that has domain $\mathcal{X}=\left\{0,1,...,n\right\}$ for any $n\in\mathbb{N}$.
An incomplete Boolean matrix is a matrix with entries in $\left\{-1,1,*\right\}$, where undefined entries are filled with ***.
A submatrix is a matrix that is obtained by extracting certain rows and/or columns from a given matrix.
A half-integer is a number of the form $n+1/2$, where $n\in\mathbb{Z}$.
We introduce some Boolean operators as follows. For every $n\in\mathbb{N}$ and $x,y\in\left\{0,1\right\}^{n}$:
  • $\overline{x}\coloneqq\left\{\overline{x_{0}},...,\overline{x_{n-1}}\right\}=\left\{1-x_{0},...,1-x_{n-1}\right\}$;
  • $x\land y\coloneqq\left\{x_{0}\land y_{0},...,x_{n-1}\land y_{n-1}\right\}$; and
  • $x\oplus y\coloneqq\left\{x_{0}\oplus y_{0},...,x_{n-1}\oplus y_{n-1}\right\}$.
2.2 Communication Complexity Model
In the two-party communication model, Alice is given input $x$, and Bob is given input $y$. Then they aim to compute $f(x,y)$ for some function $f:\{0,1\}^{n}\times\{0,1\}^{n}\to\{-1,1,*\}$ by communication protocols. The deterministic communication complexity $D(f)$ is defined as the cost of the deterministic protocol with the smallest communication cost, which computes $f$ correctly on any input. The randomized communication complexity $R_{\epsilon}(f)$ is defined as the cost of the randomized protocol with the smallest communication cost, which has access to public randomness and computes $f$ correctly on any input with probability at least $1-\epsilon$. Similarly, the quantum communication complexity $Q(f)$ is defined as the cost of the quantum protocol with the smallest cost, which is not allowed to share prior entanglement, has access to public randomness and computes $f$ correctly on any input with probability at least $1-\epsilon$. If the quantum protocol is allowed with prior entanglement initially, then the corresponding quantum communication complexity is denoted $Q^{*}(f)$. If a protocol succeeds with probability at least $1-\epsilon$ on any input for some constant $\epsilon<1/2$, we say the protocol is with bounded error. If $\epsilon=1/3$, we abbreviate $R_{\epsilon}(f),Q_{\epsilon}(f),Q_{\epsilon}^{*}(f)$ as $R(f),Q(f),Q^{*}(f)$.
2.3 Permutation-Invariant Functions
In the two-party communication model, the function value of a permutation-invariant function is invariant if we perform the same permutation to the inputs of Alice and Bob. Specifically, the formal definition is as follows.
Definition 2.1 (Permutation-invariant functions [24]).
A (total or partial) function $f:\{0,1\}^{n}\times\{0,1\}^{n}\to\{-1,1,*\}$ is permutation-invariant if for all $x,y\in\{0,1\}^{n}$, and every bijection $\pi:\left\{0,...,n-1\right\}\to\left\{0,...,n-1\right\}$, $f(x^{\pi},y^{\pi})=f(x,y)$, where $x^{\pi}$ satisfies that $x^{\pi}_{(i)}=x_{\pi(i)}$ for any $i\in\left\{0,...,n-1\right\}$.
Note that any permutation-invariant function $f$ in Definition 2.1 depends only on $|x|,|y|$ and $|x\land y|$. Here ||$\cdot$|| ⋅ | is the Hamming weight for the binary string, i.e., the number of 1’s in the string. Thus, for any $a,b\in[n]$, there exists a function $f_{a,b}\colon\left\{\max\left\{0,a+b-n\right\},...,\min\left\{a,b\right\}\right\}\rightarrow\left\{-1,1\right\}$ such that
$f_{a,b}(|x\land y|)=f(x,y),$ (1)
for any $x,y\in\left\{0,1\right\}^{n}$ satisfying $|x|=a,|y|=b$. If there exist $a,b\in[n]$ such that $f_{a,b}$ is not a constant function, we say $f$ is non-trivial.
The following definition of jumps partitions the domain of $f_{a,b}$ into different intervals according to the transition of function values.
Definition 2.2 (Jump in $f_{a,b}$ [24]).
$(c,g)$ is a jump in $f_{a,b}$ if
  1. 1.
    $f_{a,b}(c-g)\neq f_{a,b}(c+g)$;
  2. 2.
    $f_{a,b}(c-g),f_{a,b}(c+g)\in\{-1,1\}$;
  3. 3.
    $f_{a,b}(r)$ is undefined for $c-g<r<c+g$.
Moreover, we define $\mathscr{J}(f_{a,b})$ to be the set of all jumps in $f_{a,b}$:
[…]
The following definition gives an important instance of permutation-invariant functions.
Definition 2.3 (Set-Inclusion Problem).
We define the Set-Inclusion Problem $\mathsf{SetInc}^{n}_{a,b,c,g}$ as the following partial function:
$\mathsf{SetInc}_{a,b,c,g}^{n}(x,y)\coloneqq\begin{cases}-1&\text{if $|x|=a,|y|=b$ and $|x\land y|\leq c-g$},\\ 1&\text{if $|x|=a,|y|=b$ and $|x\land y|\geq c+g$},\\ *&\text{otherwise.}\end{cases}$
Additionally, we define the Exact Set-Inclusion Problem $\mathsf{ESetInc}_{a,b,c,g}^{n}$ as follows.
$\mathsf{ESetInc}_{a,b,c,g}^{n}(x,y)\coloneqq\begin{cases}-1&\text{if $|x|=a,|y|=b$ and $|x\land y|=c-g$},\\ 1&\text{if $|x|=a,|y|=b$ and $|x\land y|=c+g$},\\ *&\text{otherwise.}\end{cases}$
2.4 Rank and Approximate Rank
If $F$ is a real matrix, let $\operatorname{\mathrm{rank}}(F)$ be the rank of $F$. Then we define the approximate rank for any incomplete matrix as follows.
Definition 2.4 (Approximate rank).
For an incomplete matrix $F\in\left\{-1,1,*\right\}^{m\times n}$ and $10\leq\epsilon<$ 1, we say a real matrix $A$ approximates $F$ with error $\epsilon$ if:
  1. (1)
    $|A_{i,j}-F_{i,j}|\leq\epsilon$ for any $i\in[m],j\in[n]$ such that $F_{i,j}\neq*$;
  2. (2)
    $|A_{i,j}|\leq 1$ for all $i\in[m],j\in[n]$.
Let $\mathcal{F}_{\epsilon}$ be the set of all the real matrices that approximate $F$ with error $\epsilon$. The approximate rank of $F$ with error $\epsilon$, denoted by $\operatorname{\widetilde{rank}}_{\epsilon}(F)$, is the least rank among all real matrices in $\mathcal{F}_{\epsilon}$. If $\epsilon=2/3$, we abbreviate $\operatorname{\widetilde{rank}}_{\epsilon}(F)$ as $\operatorname{\widetilde{rank}}(F)$.
Let $f$ be a Boolean function, $\operatorname{\mathrm{rank}}(f)\coloneqq\operatorname{\mathrm{rank}}\left(M_{f}\right)$ and $\operatorname{\widetilde{rank}}(f)\coloneqq\operatorname{\widetilde{rank}}{\left(M_{f}\right)}$, where $M_{f}$ is the input matrix of $f$.
3 Polynomial Equivalence on Communication Complexity of Permutation-Invariant Functions
To show the polynomial equivalence between quantum and classical communication complexity of permutation-invariant functions as stated in Theorem 1.5, we prove the following two theorems (proved in Sections 3.1 and 3.2, respectively) for the quantum and randomized communication complexities of permutation-invariant functions using the measure in Definition 3.1.
Definition 3.1 (Measure $m(f)$).
Fix $n\in\mathbb{Z}$. Let $f:\{0,1\}^{n}\times\{0,1\}^{n}\to\{0,1,*\}$ be a permutation-invariant function. We define the measure $m(f)$ of $f$ as follows:
$m(f)\coloneqq\max_{\begin{subarray}{c}a,b\in[n]\\ (c,g)\in\mathscr{J}\left(f_{a,b}\right)\\ n_{1}\coloneqq\min\{[a-c,c,b-c,n-a-b+c]\}\\ n_{2}\coloneqq\min\left(\{[a-c,c,b-c,n-a-b+c]\}\setminus\left\{n_{1}\right\}\right)\end{subarray}}\frac{\sqrt{n_{1}n_{2}}}{g}\enspace.$
Note that this definition is motivated by Lemma 3.4.
The measure $m(\cdot)$ is inspired by the complexity measure introduced in [24], which was used to capture the randomized communication complexity of permutation-invariant functions.
Theorem 3.2 (Lower Bound).
Fix $n\in\mathbb{Z}$. Let $f:\{0,1\}^{n}\times\{0,1\}^{n}\to\{0,1,*\}$ be a permutation-invariant function. We have
$Q^{*}(f)=\Omega(m(f))\enspace.$
Theorem 3.3 (Upper Bound).
Fix $n\in\mathbb{N}$. Given a permutation-invariant function $f:\{0,1\}^{n}\times\{0,1\}^{n}\to\{0,1,*\}$ and the corresponding measure $m(f)$ defined in Definition 3.1, we have
  • $R(f)=O\left(m(f)^{2}\log^{2}n\log\log n\right)$, and
  • $Q(f)=O\left(m(f)\log^{2}n\log\log n\right)$.
3.1 Quantum Communication Complexity Lower Bound
In this section, our goal is to obtain a lower bound on the quantum communication complexity for permutation-invariant functions (Theorem 3.2). Towards this end, we show that every permutation-invariant function $f$ can be reduced to $\mathsf{ESetInc}$ (defined in Definition 2.3) and exhibit a lower bound for $\mathsf{ESetInc}$ (Lemma 3.4). Additionally, Lemma 3.4 implies if $|x|=a,|y|=b$, then the cost to distinguish $|x\land y|=c-g$ from $|x\land y|=c+g$ is related to the smallest and the second smallest number in ${[a-c,c,b-c,n-a-b+c]}$.
Lemma 3.4.
Fix $n,a,b\in\mathbb{Z}$. Consider $c$ and $g$ such that $c+g,c-g\in\mathbb{Z}$. Let $n_{1}\coloneqq\min\{[a-c,c,b-c,n-a-b+c]\}$ and $n_{2}\coloneqq\min\left(\{[a-c,c,b-c,n-a-b+c]\}\setminus\left\{n_{1}\right\}\right)$. We have
$Q^{*}\left(\mathsf{ESetInc}_{a,b,c,g}^{n}\right)=\Omega\left(\frac{\sqrt{n_{1}n_{2}}}{g}\right).$
Proof of Theorem 3.2.
By the definitions of $f_{a,b}$ and jump of $f_{a,b}$, any quantum protocol computing $f$ can also compute $\mathsf{ESetInc}^{n}_{a,b,c,g}$ for any $a,b$ and any jump $(c,g)\in\mathscr{J}(f_{a,b})$. Therefore, given a jump $(c,g)$ for $f_{a,b}$, the cost of computing $\mathsf{ESetInc}^{n}_{a,b,c,g}$ lower bounds the cost of computing $f$. By Lemma 3.4, we have $Q^{*}(f)\geq\frac{\sqrt{n_{1}n_{2}}}{g}$ for any jump $(c,g)$ in $f_{a,b}$, where $n_{1},n_{2}$ are the smallest and the second smallest number in $\{[a-c,c,b-c,n-a-b+c]\}$. We conclude that $Q^{*}(f)=\Omega\left(m(f)\right)$ as desired. ∎
Now we remain to show Lemma 3.4. We note that the following two lemmas imply Lemma 3.4 directly, where Lemma 3.5 reduces the instance such that the parameter only relies on $n_{1},n_{2},g$ and Lemma 3.6 gives the final lower bound.
Lemma 3.5.
Fix $n,a,b\in\mathbb{Z}$. Consider $c$ and $g$ such that $c+g,c-g\in\mathbb{Z}$. Let $n_{1}\coloneqq\min\{[a-c,c,b-c,n-a-b+c]\}$ and $n_{2}\coloneqq\min\left(\{[a-c,c,b-c,n-a-b+c]\}\setminus\left\{n_{1}\right\}\right)$. We have
$Q^{*}\left(\mathsf{ESetInc}_{a,b,c,g}^{n}\right)\geq Q^{*}\left(\mathsf{ESetInc}_{n_{1}+n_{2},n_{1}+n_{2},n_{1},g}^{n_{1}+3n_{2}}\right).$
Lemma 3.6.
For $n_{1},n_{2}\in\mathbb{N}$ such that $n_{1}\leq n_{2}$, we have
$Q^{*}\left(\mathsf{ESetInc}_{n_{1}+n_{2},n_{1}+n_{2},n_{1},g}^{n_{1}+3n_{2}}\right)=\Omega\left(\frac{\sqrt{n_{1}n_{2}}}{g}\right).$
We use the following two results on $\mathsf{ESetInc}$ to show Lemmas 3.5 and 3.6. Specifically, Lemma 3.7 is a variant of Lemma 4.1 in [19] and shows some reduction methods to the instances of the Exact Set-Inclusion Problem. Lemma 3.8 is a generalization of Theorem 5 in [9] proved by pattern matrix method and shows the lower bound of a special instance of the Exact Set-Inclusion Problem. The proofs of Lemmas 3.7 and 3.8 are given in Appendix B.1.
Lemma 3.7.
Fix $n,a,b\in\mathbb{Z}$. Consider $c$ and $g$ such that $c+g,c-g\in\mathbb{Z}$. The following relations hold.
  1. 1.
    $Q^{*}\left(\mathsf{ESetInc}_{a,b,c,g}^{n}\right)\leq Q^{*}\left(\mathsf{ESetInc}_{a+\ell_{1}+\ell_{3},b+\ell_{2}+\ell_{3},c+\ell_{3},g}^{n+\ell}\right)$ for integers $\ell_{1},\ell_{2},\ell_{3}\geq 0$ such that $\ell_{1}+\ell_{2}+\ell_{3}\leq\ell$;
  2. 2.
    $Q^{*}\left(\mathsf{ESetInc}_{a,b,c,g}^{n}\right)=Q^{*}\left(\mathsf{ESetInc}_{a,n-b,a-c,g}^{n}\right)=Q^{*}\left(\mathsf{ESetInc}_{n-a,b,b-c,g}^{n}\right)$;
  3. 3.
    $Q^{*}\left(\mathsf{ESetInc}_{a,b,c,g}^{n}\right)\leq Q^{*}\left(\mathsf{ESetInc}_{ka,kb,kc,kg}^{kn}\right)$, where $k\geq 1$ is an integer.
Lemma 3.8.
For every $k\in\mathbb{Z}$, if $l$ is a half-integer and $l\leq k/2$, then $Q^{*}\left(\mathsf{ESetInc}^{4k}_{2k,k,l,1/2}\right)=\Omega\left(\sqrt{kl}\right)$.
Proof of Lemma 3.5.
Using the second item of Lemma 3.7, we assume $n_{1}=c$ without loss of generality. Furthermore, we assume $n_{2}=a-c$. Let $n_{3}\coloneqq b-c,n_{4}\coloneqq n-a-b+c$. Then $n_{3},n_{4}\geq n_{2}\geq n_{1}$ and $n=n_{1}+n_{2}+n_{3}+n_{4}$. By Lemma 3.7, we have
$Q^{*}\left(\mathsf{ESetInc}^{n_{1}+3n_{2}}_{n_{1}+n_{2},n_{1}+n_{2},n_{1},g}\right)$ $=Q^{*}\left(\mathsf{ESetInc}^{n_{1}+n_{2}+n_{2}+n_{2}}_{n_{1}+n_{2},n_{1}+n_{2},n_{1},g}\right)$
$\leq Q^{*}\left(\mathsf{ESetInc}^{n_{1}+n_{2}+n_{3}+n_{4}}_{n_{1}+n_{2},n_{1}+n_{3},n_{1},g}\right)$
$=Q^{*}\left(\mathsf{ESetInc}^{n}_{a,b,c,g}\right).$
If $n_{2}=b-c$ or $n-a-b+c$, the argument is similar. ∎
Proof of Lemma 3.6.
Let $m_{1}=\left\lfloor\frac{n_{1}}{2g}+\frac{1}{2}\right\rfloor-\frac{1}{2}$, i.e., $m_{1}$ is the largest half-integer no more than $\frac{n_{1}}{2g}$. Similarly, let $m_{2}=\left\lfloor\frac{n_{2}}{2g}+\frac{1}{2}\right\rfloor-\frac{1}{2}$. By Lemma 3.7, we have
$Q^{*}\left(\mathsf{ESetInc}^{n_{1}+3n_{2}}_{n_{1}+n_{2},n_{1}+n_{2},n_{1},g}\right)\geq Q^{*}\left(\mathsf{ESetInc}^{m_{1}+3m_{2}}_{m_{1}+m_{2},m_{1}+m_{2},m_{1},1/2}\right).$
Then we discuss the following three cases:
  • Case 1: $m_{1}=m_{2}=1/2$. We have
    $Q^{*}\left(\mathsf{ESetInc}^{m_{1}+3m_{2}}_{m_{1}+m_{2},m_{1}+m_{2},m_{1},1/2}\right)=\Omega\left(1\right)=\Omega\left(\sqrt{m_{1}m_{2}}\right).$
  • Case 2: $m_{2}\geq 3/2$ and $m_{1}=1/2$. Let $m^{\prime}_{2}\coloneqq\left\lfloor\frac{m_{1}+m_{2}}{2}\right\rfloor,l_{1}\coloneqq m_{1}+m_{2}-2m^{\prime}_{2},l_{2}\coloneqq m_{1}+m_{2}-m^{\prime}_{2},l\coloneqq m_{1}+3m_{2}-4m^{\prime}_{2}$. Then,
    $l-(l_{1}+l_{2})=m_{2}-m_{1}-m^{\prime}_{2}\geq\frac{m_{2}+m_{1}}{2}-m^{\prime}_{2}\geq 0.$
    By LABEL:{le:SecInc} and 3.8, we have
    $Q^{*}\left(\mathsf{ESetInc}^{m_{1}+3m_{2}}_{m_{1}+m_{2},m_{1}+m_{2},m_{1},1/2}\right)$ $=Q^{*}\left(\mathsf{ESetInc}^{4m^{\prime}_{2}+l}_{2m^{\prime}_{2}+l_{1},m^{\prime}_{2}+l_{2},m_{1},1/2}\right)$
    $\geq Q^{*}\left(\mathsf{ESetInc}^{4m^{\prime}_{2}}_{2m^{\prime}_{2},m^{\prime}_{2},m_{1},1/2}\right)$
    $=\Omega\left(\sqrt{m_{1}m^{\prime}_{2}}\right)$
    $=\Omega\left(\sqrt{m_{1}m_{2}}\right).$
  • Case 3: $m_{1}\geq 3/2$. Let $m\coloneqq\left\lfloor\frac{m_{1}}{6}+\frac{m_{2}}{2}\right\rfloor,k\coloneqq\left\lfloor\frac{m_{1}}{3}+\frac{1}{2}\right\rfloor-\frac{1}{2},l_{3}\coloneqq m_{1}-k,l_{1}\coloneqq\left(m_{1}+m_{2}-2m\right)-l_{3},l_{2}\coloneqq\left(m_{1}+m_{2}-m\right)-l_{3},l\coloneqq m_{1}+3m_{2}-4m$. Since $k$ is the largest half-integer smaller than $\frac{m_{1}}{3}$, we have $k\leq\frac{1}{2}\cdot\left\lfloor\frac{2m_{1}}{3}\right\rfloor$. Since $m_{1}\leq m_{2}$, we have
    $k\leq\frac{1}{2}\cdot\left\lfloor\frac{2m_{1}}{3}\right\rfloor\leq\frac{1}{2}\cdot\left\lfloor\frac{m_{1}}{6}+\frac{m_{2}}{2}\right\rfloor\leq\frac{m}{2}\enspace,$ (2)
    and
    $l-\left(l_{1}+l_{2}+l_{3}\right)=m_{2}-k-m\geq m_{2}-\frac{m_{1}}{3}-\left(\frac{m_{1}}{6}+\frac{m_{2}}{2}\right)\geq 0\enspace.$ (3)
    Then we have
    $Q^{*}\left(\mathsf{ESetInc}^{m_{1}+3m_{2}}_{m_{1}+m_{2},m_{1}+m_{2},m_{1},1/2}\right)$
    =$\quad$= $Q^{*}\left(\mathsf{ESetInc}^{4m+l}_{2m+l_{1}+l_{3},m+l_{2}+l_{3},k+l_{3},1/2}\right)$
[…]
    =$\quad$= $\Omega\left(\sqrt{m_{1}m_{2}}\right)\enspace.$
We conclude that
$Q^{*}\left(\mathsf{ESetInc}^{n_{1}+3n_{2}}_{n_{1}+n_{2},n_{1}+n_{2},n_{1},g}\right)\geq Q^{*}\left(\mathsf{ESetInc}^{m_{1}+3m_{2}}_{m_{1}+m_{2},m_{1}+m_{2},m_{1},1/2}\right)=\Omega\left(\frac{\sqrt{n_{1}n_{2}}}{g}\right)\enspace.$
3.2 Randomized and Quantum Communication Complexity Upper Bound
We show upper bounds on the randomized and quantum communication complexities for permutation invariant functions (Theorem 3.3). Similar to Section 3.1, we do so by giving upper bounds for a specific problem, $\mathsf{SetInc}$ (see Definition 2.3), and reducing permutation-invariant functions to $\mathsf{SetInc}$.
The following two lemmas capture the randomized and quantum communication complexity for $\mathsf{SetInc}$, respectively.
Lemma 3.9 (Classical Upper Bound).
Fix $n,a,b\in\mathbb{Z}$. Consider $c,g$ such that $c+g,c-g\in\mathbb{Z}$. Let $n_{1}\coloneqq\min\{[a-c,c,b-c,n-a-b+c]\}$ and $n_{2}\coloneqq\min\left(\{[a-c,c,b-c,n-a-b+c]\}\setminus\left\{n_{1}\right\}\right)$. For any input $x,y\in\{0,1\}^{n}$ of $\mathsf{SetInc}_{a,b,c,g}^{n}$, there exists a randomized communication protocol that computes $\mathsf{SetInc}_{a,b,c,g}^{n}(x,y)$ using $O\left(\frac{n_{1}n_{2}}{g^{2}}\log n\log\log n\right)$ bits of communication with success probability at least $1-1/(6\log n)$.
Lemma 3.10 (Quantum Upper Bound).
Fix $n,a,b\in\mathbb{Z}$. Consider $c,g$ such that $c+g,c-g\in\mathbb{Z}$. Let $n_{1}\coloneqq\min\{[a-c,c,b-c,n-a-b+c]\}$ and $n_{2}\coloneqq\min\left(\{[a-c,c,b-c,n-a-b+c]\}\setminus\left\{n_{1}\right\}\right)$. For any input $x,y\in\{0,1\}^{n}$ of $\mathsf{SetInc}_{a,b,c,g}^{n}$, there exists a quantum communication protocol without prior entanglement that computes $\mathsf{SetInc}_{a,b,c,g}^{n}(x,y)$ using $O\left(\frac{\sqrt{n_{1}n_{2}}}{g}\log n\log\log n\right)$ qubits of communication with success probability at least $1-1/(6\log n)$.
We note that Lemma 3.10 is a quantum speedup version of Lemma 3.9 by quantum amplitude amplification. The proof of Lemma 3.9 is given at the end of this section, and we postpone the proof of Lemma 3.10 to Section B.2.
Now we explain how to derive Theorem 3.3 from the lemmas above.
Proof of Theorem 3.3.
We first present a randomized protocol to compute $f$ based on binary search:
  1. 1.
    Alice and Bob exchange $a\coloneqq|x|,b\coloneqq|y|$.
  2. 2.
    Alice and Bob both derive $f_{a,b}$ such that $f_{a,b}(|x\land y|)=f(x,y)$.
  3. 3.
    Let $\mathscr{J}\left(f_{a,b}\right)=\{(c_{i},g_{i})\}_{i\in[k]}$ for some $k\leq n$ be the set of jumps of $f_{a,b}$ as in Definition 2.2.
  4. 4.
    Alice and Bob use binary search to determine $i\in\{0,1,...,k\}$ such that $\lvert x\land y\rvert\in I_{i}$, where $I_{i}$ is defined in Equation (4).
We first discuss the communication complexity of the above protocol. The first step takes $O(\log n)$ bits of communication. The fourth step costs $O\left(m(f)^{2}\log^{2}n\log\log n\right)$ bits of communication: For each $i\in[k]$, Alice and Bob can determine whether $|x\land y|\leq c_{i}-g_{i}$ or $|x\land y|\geq c_{i}+g_{i}$ by $O(m(f)^{2}\log n\log\log n)$ with a success probability of at least $1-1/\left(6\log n\right)$ by Lemma 3.9. Since binary search takes at most $\lceil\log\left(k+1\right)\rceil=O\left(\log n\right)$ rounds, the total communication cost is $O\left(m(f)^{2}\log^{2}n\log\log n\right)$.
Now we argue for the correctness of the protocol. Notice that the set of jumps $\mathscr{J}\left(f_{a,b}\right)$ invokes $k+1$ intervals:
$\left\{I_{0}\coloneqq[0,c_{1}-g_{1}],I_{1}\coloneqq[c_{1}+g_{1},c_{2}-g_{2}],\dots,I_{k-1}\coloneqq[c_{k-1}+g_{k-1},c_{k}-g_{k}],I_{k}\coloneqq[c_{k}+g_{k},n]\right\}\enspace.$ (4)
In particular, the followings hold:
  • For every $j\in[0,k]$ and $z_{1},z_{2}\in I_{j}$ such that $f_{a,b}(z_{1})\neq*$ and $f_{a,b}(z_{2})\neq*$, we have $f_{a,b}(z_{1})=f_{a,b}(z_{2})$.
  • If $z\notin I_{j}$ for any $j\in[0,k]$, then $f_{a,b}(z)=*$.
Therefore, Alice and Bob start from $i=\lfloor(k+1)/2\rfloor$ to determine whether $|x\land y|\leq c_{i}-g_{i}$ or $|x\land y|\geq c_{i}+g_{i}$ with success probability of at least $1-1/\left(6\log n\right)$. Depending on the result, they repeat the same process similar to binary search to find the interval that $|x\land y|$ falls in. After at most $\lceil\log\left(k+1\right)\rceil=O\left(\log n\right)$ repetitions, there is only one remaining interval and they can easily determine $f_{a,b}(|x\land y|)$. For $n\geq 2$, the failure probability of the above protocol is at most
$1-\left(1-\frac{1}{6\log n}\right)^{\lceil\log\left(k+1\right)\rceil}\leq\frac{\lceil\log\left(k+1\right)\rceil}{6\log n}\leq\frac{\lceil\log\left(n+1\right)\rceil}{3\log n^{2}}\leq\frac{1}{3}.$
For the quantum case, Alice and Bob use the same protocol above, but we invoke Lemma 3.10 to analyze the communication complexity.
Proof of Lemma 3.9.
We rely on the following two claims to prove the lemma.
Fact 3.11 ([2, Lemma 30]).
Fix $10<\epsilon<\beta<$ 1 such that $\beta+\epsilon\leq 1$. For a set $S$, suppose there is a subset $S^{\prime}$ of $S$ such that $\frac{|S^{\prime}|}{|S|}\leq\beta-\epsilon$ or $\frac{|S^{\prime}|}{|S|}\geq\beta+\epsilon$. Suppose we can sample from $S$ uniformly and ask whether the sample is in $S^{\prime}$. Then we can decide whether $\frac{|S^{\prime}|}{|S|}\leq\beta-\epsilon$ or $\frac{|S^{\prime}|}{|S|}\geq\beta+\epsilon$ by $O(\beta/\epsilon^{2})$ samples, with success probability at least $2/3$.
Fact 3.12.
Suppose $x,y\in\{0,1\}^{n}$ are the inputs of Alice and Bob such that $|x|\neq|y|$. Alice and Bob can sample an element from $S\coloneqq\left\{i:x_{i}\neq y_{i}\right\}$ uniformly using $O(\log n)$ bits of communication.
We refer interesting readers to Section A.4 for the proof of 3.12. Now we prove the lemma by casing on the values of $n_{1}$ and $n_{2}$.
  • Case 1: $n_{1}=c$ and $n_{2}=a-c$. According to Definition 2.3, we have either $\frac{|x\land y|}{|x|}\leq\frac{c-g}{a}$ or $\frac{|x\land y|}{|x|}\geq\frac{c+g}{a}$. Alice and Bob estimate $\frac{|x\land y|}{|x|}$ as follows: Alice chooses an index $i$ such that $x_{i}=1$ uniformly at random. Then Alice sends $i$ to Bob, and Bob checks whether $y_{i}=1$. By 3.11, setting $\beta\coloneqq\frac{c}{a}$, $\epsilon\coloneqq\frac{g}{a}$, Alice and Bob can decide whether $\frac{|x\land y|}{|x|}\leq\frac{c-g}{a}$ or $\frac{|x\land y|}{|x|}\geq\frac{c+g}{a}$ with bounded error using $O\left(\frac{ac}{g^{2}}\right)=O\left(\frac{n_{1}n_{2}}{g^{2}}\right)$ samples. Since $|x|=a$, using $O\left(\frac{n_{1}n_{2}}{g^{2}}\log\log n\right)$ samples, they can decide whether $|x\land y|\leq c-g$ or $|x\land y|\geq c+g$ with success probability at least $1-1/(6\log n)$ by error reduction. Thus, the communication complexity is $O\left(\frac{n_{1}n_{2}}{g^{2}}\log n\log\log n\right)$.
  • Case 2: $n_{1}=a-c$ and $n_{2}=c$, or $n_{1}=a-c$ and $n_{2}=c$, or $n_{1}=c$ and $n_{2}=b-c$. A similar argument as in Case 1 applies.
  • Case 3: $n_{1}=c$ and $n_{2}=n-a-b+c$. Since $n_{1}\leq n_{2}$, we have $a+b\leq n$. Then we consider the following two cases:
    1. 1.
      Case 3.1: $a+b<n$. Let $m\coloneqq n_{1}+n_{2}$, $p\coloneqq\frac{|x\land y|}{|\overline{x}\oplus y|}$. Since
[…]
      we have
      $p=\frac{|x\land y|}{n-(a+b)+2|x\land y|}=\frac{1}{\frac{n-(a+b)}{|x\land y|}+2}\enspace.$
      Notice that $p$ is an increasing function with respect to $|x\land y|$. As a result, if $|x\land y|\leq c-g$, then $p\leq\frac{c-g}{m-2g}$; if $|x\land y|\geq c+g$, then $p\geq\frac{c+g}{m+2g}$. Let
      $\beta\coloneqq\frac{1}{2}\left(\frac{c+g}{m+2g}+\frac{c-g}{m-2g}\right)=\frac{cm-2g^{2}}{m^{2}-4g^{2}}\text{ and }\epsilon\coloneqq\frac{1}{2}\left(\frac{c+g}{m+2g}-\frac{c-g}{m-2g}\right)=\frac{gm}{m^{2}-4g^{2}}\enspace.$
      Since $c-g\geq 0$, we have $\beta\leq\frac{c+g}{m+2g}\leq\frac{2c}{m}$ and
      $\epsilon$ $=\frac{1}{2}\left(\frac{c+g}{m+2g}-\frac{c}{m}\right)+\frac{1}{2}\left(\frac{c}{m}-\frac{c-g}{m-2g}\right)$
      $=\frac{1}{2}\left(\frac{g(m-2c)}{m(m+2g)}+\frac{g(m-2c)}{m(m-2g)}\right)$
      $=O\left(\frac{g}{m}\right)\enspace.$
      For any $x\in\{0,1\}^{n}$, we let $S_{x}\coloneqq\left\{i:x_{i}=1\right\}$. By 3.12, Alice and Bob can sample $i$ from $S_{\overline{x}\oplus y}$ uniformly using $O(\log n)$ bits communication. Since $i\in S_{\overline{x}\oplus y}$, if $x_{i}=y_{i}=1$, then $i\in S_{x\land y}$; if $x_{i}=y_{i}=0$, then $i\notin S_{x\land y}$. By 3.11, using $O\left(\frac{\beta}{\epsilon^{2}}\right)=O\left(\frac{mc}{g^{2}}\right)=O\left(\frac{n_{1}n_{2}}{g^{2}}\right)$ samples, Alice and Bob can decide whether $p\leq\beta-\epsilon$ or $p\geq\beta+\epsilon$ with bounded error. Equivalently, Alice and Bob can distinguish $|x\land y|\leq c-g$ from $|x\land y|\geq c+g$ with bounded error. By error reduction, using $O\left(\frac{n_{1}n_{2}}{g^{2}}\log\log n\right)$ samples, they can decide whether $|x\land y|\geq c-g$ or $|x\land y|\leq c+g$ with success probability at least $1-1/(6\log n)$. Thus, the communication complexity is $O\left(\frac{n_{1}n_{2}}{g^{2}}\log n\log\log n\right)$.
    2. 2.
      Case 3.2: $a+b=n$. Alice and Bob generate new inputs $x^{\prime}=x0$ and $y^{\prime}=y0$ (pad a zero after the original input). We know
      $\mathsf{SetInc}_{a,b,c,g}^{n}(x,y)=\mathsf{SetInc}_{a,b,c,g}^{n+1}(x^{\prime},y^{\prime})\enspace.$
      Since $a+b<n+1$, Alice and Bob perform the protocol in Case 3.1 in the new inputs, and the complexity analysis is similar to Case 3.1.
  • Case 4: $n_{1}=n-a-b+c$ and $n_{2}=c$, or $n_{1}=a-c$ and $n_{2}=b-c$, or $n_{1}=b-c$ and $n_{2}=a-c$. A similar argument as in Case 3 works.
4 Log-Rank Conjecture for Permutation-Invariant Functions
Theorem 1.6 states the Log-rank Conjecture for permutation-invariants functions. We argue for the lower bound (Lemma 4.1) and the upper bound (Lemma 4.2) separately.
Lemma 4.1.
Fix $n\in\mathbb{N}$. Let $f:\{0,1\}^{n}\times\{0,1\}^{n}\rightarrow\left\{-1,1\right\}$ be a non-trivial total permutation-invariant function. For every $a,b\in[n]$ such that $f_{a,b}$ is not a constant function, we have
$\log\operatorname{\mathrm{rank}}(f)=\Omega\left(\max\left\{\log n,\min\left\{a,b,n-a,n-b\right\}\right\}\right),$
where $f_{a,b}$ is defined as Equation 1.
Lemma 4.2.
Fix $n\in\mathbb{N}$. Let $f:\{0,1\}^{n}\times\{0,1\}^{n}\rightarrow\left\{-1,1\right\}$ be a non-trivial total permutation-invariant function.
$D(f)=O\left(\max_{a,b\in[n]:f_{a,b}\text{ is not constant}}\min\left\{a,b,n-a,n-b\right\}\cdot\log n\right)\enspace,$
where $f_{a,b}$ is defined as Equation 1.
We prove Lemma 4.1 below, and postpone the proof of Lemma 4.2 to Appendix C.
Proof of Lemma 4.1.
We rely on the following two claims to prove the lemma. Two claims show the lower bound on the rank of some special functions respectively.
Fact 4.3 ([16], merging Corollary 6 with Lemma 4).
Fix $n\in\mathbb{N}$. Let $f:\{0,1\}^{n}\times\{0,1\}^{n}\rightarrow\left\{-1,1\right\}$ be defined as $f(x,y)\coloneqq D(|x\land y|)$ for some predicate $D:\left\{0,1,...,n\right\}\rightarrow\left\{-1,1\right\}$. If $t$ is the smallest integer such that $D(t)\neq D(t-1)$, then $\log\operatorname{\mathrm{rank}}(f)=\Omega\left(\log\left(\sum_{i=t}^{n}\binom{n}{i}\right)\right)$.
Fact 4.4.
Fix $n\in\mathbb{N}$. Let $\mathcal{X},\mathcal{Y}\coloneqq\left\{x\in\{0,1\}^{n}:|x|=k\right\}$, where $k\leq n/2$. Let $\operatorname{\mathsf{DISJ}}_{n}^{k}:\mathcal{X}\times\mathcal{Y}\rightarrow\left\{-1,1\right\}$ and $\operatorname{\mathsf{EQ}}_{n}^{k}:\mathcal{X}\times\mathcal{Y}\rightarrow\left\{-1,1\right\}$ be defined as
[…]
Then $\operatorname{\mathrm{rank}}\left(\operatorname{\mathsf{DISJ}}_{n}^{k}\right)\geq\binom{n}{k}-1$ and $\operatorname{\mathrm{rank}}\left(\operatorname{\mathsf{EQ}}_{n}^{k}\right)\geq\binom{n}{k}-1$.
We refer interesting readers to Section A.2 for the proof of 4.4. Now we prove the lemma by casing on the values of $a$ and $b$.
We can assume $a\leq b\leq n/2$ without loss of generality because the cases where $a>n/2$ or $b>n/2$ can be obtained by flipping each bit of Alice or Bob’s input. Thus, it suffices to prove $\log\operatorname{\mathrm{rank}}(f)=\Omega\left(\max\left\{\log n,a\right\}\right)$.
We prove the following two claims that directly lead to our result:
  1. 1.
    If $a\leq b\leq n/2$ and $a=o\left(\log n\right)$, then $\log\operatorname{\mathrm{rank}}(f)=\Omega\left(\log n\right)$.
  2. 2.
    If $a\leq b\leq n/2$ and $a=\Omega\left(\log n\right)$, then $\log\operatorname{\mathrm{rank}}(f)=\Omega\left(a\right)$.
We first prove Item 1. Suppose $a\leq b\leq n/2$ and $a=o\left(\log n\right)$. Since $f_{a,b}$ is not a constant function, there exists $c\in[0,a)$ such that $f_{a,b}(c)\neq f_{a,b}(c+1)$. Without loss of generality, we assume $f_{a,b}(c)=-1$. Let $n^{\prime}\coloneqq n-(a+b-c-2)$. Since $b\leq n/2$ and $c\leq a=o(\log n)$, $n^{\prime}=n-(a+b-c-2)=\Omega\left(n\right)$. Let $\mathcal{X}$ and $\mathcal{Y}$ be the set $\left\{x\in\{0,1\}^{n^{\prime}}:|x|=1\right\}$. For any $x\in\mathcal{X},y\in\mathcal{Y}$,
$\operatorname{\mathsf{DISJ}}_{n^{\prime}}^{1}(x,y)=f_{a,b}(|x\land y|+c)=f(x^{\prime},y^{\prime})\enspace,$
where
$x^{\prime}\coloneqq x\underbrace{1\cdots 1}_{c}\underbrace{1\cdots 1}_{a-c-1}\underbrace{0\cdots 0}_{b-c-1}\text{ and }y^{\prime}\coloneqq y\underbrace{1\cdots 1}_{c}\underbrace{0\cdots 0}_{a-c-1}\underbrace{1\cdots 1}_{b-c-1}\enspace.$
Thus, $\operatorname{\mathsf{DISJ}}_{n^{\prime}}^{1}$ is a submatrix of $f$. By 4.4, we have
$\log\operatorname{\mathrm{rank}}(f)\geq\log\operatorname{\mathrm{rank}}(\operatorname{\mathsf{DISJ}}_{n^{\prime}}^{1})\geq\log\left(n^{\prime}-1\right)=\Omega\left(\log n\right).$
Now we prove Item 2. Suppose $a,b\leq n/2$ and $\min\left\{a,b\right\}=\Omega\left(\log n\right)$, we consider the following three cases:
  • Case 1: There exists $c\in[4a/7,3a/5)$ such that $f_{a,b}(c)\neq f_{a,b}(c+1)$. Let $k=\lfloor a/2\rfloor$ and $k^{\prime}=\lceil a/2\rceil$. Let $g:\{0,1\}^{k}\times\{0,1\}^{k}\rightarrow\left\{-1,1\right\}$ be such that $g(x,y)=f_{a,b}(|x^{\prime}\land y^{\prime}|)$ for every $x,y\in\{0,1\}^{k}$, where
    $x^{\prime}\coloneqq x\overline{x}\underbrace{0\cdots 0}_{k}\underbrace{1\cdots 1}_{k^{\prime}}\underbrace{0\cdots 0}_{b-a}\underbrace{0\cdots 0}_{n-b-2k}\text{ and }y^{\prime}\coloneqq y\underbrace{0\cdots 0}_{k}\overline{y}\underbrace{1\cdots 1}_{k^{\prime}}\underbrace{1\cdots 1}_{b-a}\underbrace{0\cdots 0}_{n-b-2k}\enspace.$
    Observe that $x^{\prime},y^{\prime}\in\{0,1\}^{n}$ and $|x^{\prime}|=a,|y^{\prime}|=b$. Moreover, $g(x,y)=D(|x\land y|)$ for predicate $D:\left\{0,1,...,k\right\}\rightarrow\left\{-1,1\right\}$ such that $D(z)=f_{a,b}(z+k^{\prime})$ for every $z\in[0,k]$. Thus, we have $D(c-k^{\prime})\neq D(c-k^{\prime}+1)$. By 4.3, we have
    $\log\operatorname{\mathrm{rank}}(g)=\Omega\left(\log\left(\sum_{i=c-k^{\prime}+1}^{k}\binom{k}{i}\right)\right)\enspace.$
    Since $c-k^{\prime}+1<3a/5-\lceil a/2\rceil+1\leq a/10\leq k/2$, we conclude $\log\operatorname{\mathrm{rank}}(g)=\Omega(k)=\Omega(a)$.
  • Case 2: There exists $c\in[0,4a/7)$ such that $f_{a,b}(c)\neq f_{a,b}(c+1)$ and $f_{a,b}$ is a constant function in the range $[c,3a/5)$. Without loss of generality, we assume $f_{a,b}(c)=-1$. Let $l\coloneqq\lfloor 3a/5\rfloor$, $l^{\prime}\coloneqq\lceil 2a/5\rfloor$, $m\coloneqq n-(c+b-a+2l^{\prime})$. Since $a\leq b\leq n/2$ and $c<4a/7$, we have
    $m=n-(c+b-a+2l^{\prime})\geq 2a-2l^{\prime}-c=2l-c\geq 2(l-c)\enspace.$
    Let $\mathcal{X}$ and $\mathcal{Y}$ be the set $\left\{x\in\{0,1\}^{m}:|x|=l-c\right\}$. For every $x\in\mathcal{X},y\in\mathcal{Y}$, we have
    $\operatorname{\mathsf{DISJ}}_{m}^{l-c}(x,y)=f_{a,b}(|x^{\prime}\land y^{\prime}|)=f(x^{\prime},y^{\prime})\enspace,$
    where
    $x^{\prime}\coloneqq x\underbrace{1\cdots 1}_{c}\underbrace{0\cdots 0}_{b-a}\underbrace{0\cdots 0}_{l^{\prime}}\underbrace{1\cdots 1}_{l^{\prime}}\text{ and }y^{\prime}\coloneqq y\underbrace{1\cdots 1}_{c}\underbrace{1\cdots 1}_{b-a}\underbrace{1\cdots 1}_{l^{\prime}}\underbrace{0\cdots 0}_{l^{\prime}}.$
    Thus, $\operatorname{\mathsf{DISJ}}_{m}^{l-c}$ is a submatrix of $f$. By 4.4, we have
    $\log\operatorname{\mathrm{rank}}\left(f\right)\geq\log\operatorname{\mathrm{rank}}(\operatorname{\mathsf{DISJ}}_{m}^{l-c})=\Omega\left(\log\binom{m}{l-c}\right)=\Omega\left(l-c\right)=\Omega\left(a\right)\enspace.$
  • Case 3: There exists $c\in[3a/5,a)$ such that $f_{a,b}(c)\neq f_{a,b}(c+1)$ and $f_{a,b}$ is a constant function in the range $[0,c)$. Without loss of generality, we assume $f_{a,b}(c)=-1$. Since $a\leq b\leq\frac{n}{2}$, we have $n-b+c\geq a+c\geq 2c$. Let $\mathcal{X}$ and $\mathcal{Y}$ be the set $\left\{x\in\{0,1\}^{n-b+c}:|x|=c\right\}$. For every $x\in\mathcal{X},y\in\mathcal{Y}$, we have
    $\operatorname{\mathsf{EQ}}_{n-b+c}^{c}(x,y)=f_{a,b}(|x^{\prime}\land y^{\prime}|)=f(x^{\prime},y^{\prime}),$
    where
    $x^{\prime}\coloneqq x\underbrace{0\cdots 0}_{b-a}\underbrace{0\cdots 0}_{a-c}\text{ and }y^{\prime}\coloneqq y\underbrace{1\cdots 1}_{b-a}\underbrace{0\cdots 0}_{a-c}\enspace.$
    Thus, $\operatorname{\mathsf{EQ}}_{n-b+c}^{c}$ is a submatrix of $f$. By 4.4, we have
    $\log\operatorname{\mathrm{rank}}\left(f\right)\geq\log\operatorname{\mathrm{rank}}(\operatorname{\mathsf{EQ}}_{n-b+c}^{c})=\Omega\left(\log\binom{n-b+c}{c}\right)=\Omega\left(c\right)=\Omega\left(a\right)\enspace.$
5 Log-Approximate-Rank Conjecture for Permutation-Invariant Functions
We discuss Theorem 1.7. In particular, we use the following two lemmas (proved in Appendix D) to prove Theorem 1.7. Additionally, we note that Lemmas 5.1 and 5.2 are variants of Lemmas 3.7 and 3.8.
Lemma 5.1.
Let $n,a,b,c,g\in\mathbb{Z}^{+}$. The following relations hold:
  1. 1.
    $\operatorname{\widetilde{rank}}\left(\mathsf{ESetInc}_{a,b,c,g}^{n}\right)\leq\operatorname{\widetilde{rank}}\left(\mathsf{ESetInc}_{a+\ell_{1}+\ell_{3},b+\ell_{2}+\ell_{3},c+\ell_{3},g}^{n+\ell}\right)$ for $\ell_{1},\ell_{2},\ell_{3}\geq 0$ such that $\ell_{1}+\ell_{2}+\ell_{3}\leq\ell$;
  2. 2.
    $\operatorname{\widetilde{rank}}\left(\mathsf{ESetInc}_{a,b,c,g}^{n}\right)=\operatorname{\widetilde{rank}}\left(\mathsf{ESetInc}_{a,n-b,a-c,g}^{n}\right)=\operatorname{\widetilde{rank}}\left(\mathsf{ESetInc}_{n-a,b,b-c,g}^{n}\right)$; and
  3. 3.
    $\operatorname{\widetilde{rank}}\left(\mathsf{ESetInc}_{a,b,c,g}^{n}\right)\leq\operatorname{\widetilde{rank}}\left(\mathsf{ESetInc}_{ka,kb,kc,kg}^{kn}\right)$ for $k\geq 1$.
Lemma 5.2.
Fix $k\in\mathbb{Z}$. Let $l$ be a half-integer such that $l\leq k/2$. We have
$\log\left(\operatorname{\widetilde{rank}}\left(\mathsf{ESetInc}^{4k}_{2k,k,l,1/2}\right)\right)=\Omega\left(\sqrt{kl}\right).$
Proof sketch of Theorem 1.7.
We use a similar argument as in the proof of Lemma 3.4. Namely, for every $a,b\in[n]$ and jump $(c,g)\in\mathscr{J}(f_{a,b})$, let $n_{1}\coloneqq\min\{[a-c,c,b-c,n-a-b+c]\}$ and $n_{2}\coloneqq\min\left(\{[a-c,c,b-c,n-a-b+c]\}\setminus\left\{n_{1}\right\}\right)$. We have
$\log\operatorname{\widetilde{rank}}\left(\mathsf{ESetInc}_{a,b,c,g}^{n}\right)=\Omega\left(\frac{\sqrt{n_{1}n_{2}}}{g}\right)\enspace.$
Since $\mathsf{ESetInc}_{a,b,c,g}^{n}$ is a subfunction of $f$, we have
$\log\operatorname{\widetilde{rank}}\left(f\right)=\Omega\left(\max_{\begin{subarray}{c}a,b\in[n]\\ (c,g)\in\mathscr{J}\left(f_{a,b}\right)\end{subarray}}\frac{\sqrt{n_{1}n_{2}}}{g}\right)=\Omega\left(m(f)\right).$
Combining Theorem 3.3 and the above equation, we have Theorem 1.7 as desired. ∎
6 Conclusion
This paper proves that the randomized communication complexity of permutation-invariant Boolean functions is at most quadratic of the quantum communication complexity (up to a logarithmic factor). Our results suggest that symmetries prevent exponential quantum speedups in communication complexity, extending the analogous research on query complexity. Furthermore, we prove that the Log-rank Conjecture and Log-approximate-rank Conjecture hold for non-trivial permutation-invariant Boolean functions (up to a logarithmic factor). There are some interesting problems to explore in the future.
  • Permutation invariance over higher alphabets. In this paper, the permutation-invariant function is a binary function. The interesting question is to generalize our results to larger alphabets, i.e., to permutation-invariant functions of the form $f:{X}^{n}\times{Y}^{n}\to R$ where ${X},{Y}$ and $R$ are not necessarily binary sets.
  • Generalized permutation invariance. It is possible to generalize our results for a larger class of symmetric functions. One candidate might be a class of functions that have graph-symmetric properties. Suppose $\mathcal{G}_{A},\mathcal{G}_{B}$ are two sets of $n$-vertices graphs, and $G_{n}$ is a group that acts on the edges of an $n$-vertices graph and permutes them in a way that corresponds to relabeling the vertices of the underlying graph. A function $f:\mathcal{G}_{A}\times\mathcal{G}_{B}\to\{0,1\}$ is graph-symmetric if $f(x,y)=f(x\circ\pi,y\circ\pi)$, where $x\in\mathcal{G}_{A},y\in\mathcal{G}_{B}$, and $\pi\in G_{n}$.
Appendices Organization
The appendices are organized as follows. In Section A, we give extended preliminaries. Moreover, the relation between the sections of appendices and the omitted proofs are given in Table 1.
Table 1: The list of omitted proofs.
Section Omitted Proofs
A Facts 3.12 and 4.4 in Section 2
B Lemmas 3.7, 3.8 and 3.10 in Section 3
C Lemma 4.2 in Section 4
D Lemmas 5.1 and 5.2 in Section 5
Appendix A Extended Preliminaries
A.1 Pattern Matrix Method and Approximate Degree
Pattern matrix method [40] is a well-known method for lower bound analysis in quantum communication complexity. Fix $k,n\in\mathbb{Z}$ and $k$ divides $n$. The set $[n]$ is partitioned into $k$ blocks, each consisting of $n/k$ elements. For the universal set $[n]$, let $\mathcal{V}(n,k)$ be a family of subsets that have exactly one element in each block. Clearly, $|\mathcal{V}(n,k)|=(n/k)^{k}$. Fix $x\in\{0,1\}^{n}$ and $V\in\mathcal{V}(n,k)$, let the projection of $x$ onto $V$ be defined as $x|_{V}\coloneqq(x_{i_{1}},x_{i_{2}},\dots,x_{i_{k}})\in\{0,1\}^{k}$, where $i_{1}<i_{2}<\cdots<i_{k}$ are the elements of $V$.
Definition A.1 (Pattern matrix [40]).
The $(n,k,f)$-pattern matrix $A$ for a function $f\colon\{0,1\}^{k}\to\{-1,1,*\}$ is defined as
$A\coloneqq\Big{[}f(x|_{V}\oplus w)\Big{]}_{x\in\{0,1\}^{n},\,(V,w)\in\mathcal{V}(n,k)\times\{0,1\}^{k}}\;.$
Here, $A$ is a matrix of size $2^{n}$ by $(n/k)^{k}2^{k}$, where each row is indexed by strings $x\in\{0,1\}^{n}$ and each column is indexed by pairs $(V,w)\in\mathcal{V}(n,k)\times\{0,1\}^{k}$. The entries of $A$ are given by $A_{x,(V,w)}=f(x|_{V}\oplus w)$.
Definition A.2 (Approximate degree).
For $f:\{0,1\}^{n}\rightarrow\left\{-1,1,*\right\}$ and $10\leq\epsilon<$ 1, we say a real multilinear polynomial $p$ approximates $f$ with error $\epsilon$ if:
  1. (1)
    $|p(x)-f(x)|\leq\epsilon$ for any $x\in\{0,1\}^{n}$ such that $f(x)\neq*$;
  2. (2)
    $|p(x)|\leq 1$ for all $x\in\{0,1\}^{n}$.
The approximate degree of $f$ with error $\epsilon$, denoted by $\widetilde{\deg}_{\epsilon}(f)$, is the minimum degree among all real multilinear polynomials that approximate $f$ with error $\epsilon$. If $\epsilon=2/3$, we abbreviate $\widetilde{\deg}_{\epsilon}(f)$ as $\widetilde{\deg}(f)$.
Fact A.3 ([37], Theorem 4).
For symmetric Boolean functions $f:\{0,1\}^{n}\rightarrow\left\{-1,1,*\right\}$, let $D$ be a Boolean predicate such that $D(|x|)=f(x)$ for any $x\in\{0,1\}^{n}$. Then $\operatorname{\mathsf{\widetilde{deg}}}\left(f\right)=\Omega\left(\sqrt{n(n-\Gamma(D))}\right)$, where
$\Gamma(D)=\min\left\{|2k-n+1|:D(k),D(k+1)\neq*,D(k)\neq D(k+1)\text{ and }0\leq k\leq n-1\right\}.$
Each two-party communication problem $F:\mathcal{X}\times\mathcal{Y}\rightarrow\left\{-1,1,*\right\}$ can be viewed as a matrix naturally according to the definition of $F$. Then the following fact gives a lower bound on the quantum communication complexity by the approximate degree.
Fact A.4 ([9], Theorem 6).
Let $F$ be the $(n,t,f)$-pattern matrix, where $f\colon\{0,1\}^{t}\to\{-1,+1,*\}$ is given. Then for every $\epsilon\in[0,1)$ and every $\delta<\epsilon/2,$
$Q^{*}_{\delta}(F)$ $\geq\frac{1}{4}\operatorname{\mathsf{\widetilde{deg}}}_{\epsilon}(f)\log\left(\frac{n}{t}\right)-\frac{1}{2}\log\left(\frac{3}{\epsilon-2\delta}\right).$
A.2 Rank
We first restate and prove Fact 4.4 as follows: See 4.4
Proof.
Let ${\operatorname{\mathsf{DISJ}}^{\prime}}_{n}^{k}:\mathcal{X}\times\mathcal{Y}\rightarrow\left\{0,1\right\}$ and ${\operatorname{\mathsf{EQ}}^{\prime}}_{n}^{k}:\mathcal{X}\times\mathcal{Y}\rightarrow\left\{0,1\right\}$ be defined as
${\operatorname{\mathsf{DISJ}}^{\prime}}_{n}^{k}(x,y)\coloneqq\begin{cases}1,&\text{ if }|x\land y|=0,\\ 0,&\text{ if }|x\land y|\neq 0.\end{cases}$
and
${\operatorname{\mathsf{EQ}}^{\prime}}_{n}^{k}(x,y)\coloneqq\begin{cases}1,&\text{ if }x=y,\\ 0,&\text{ if }x\neq y.\end{cases}$
By Example 2.12 in [32], $\operatorname{\mathrm{rank}}\left({\operatorname{\mathsf{DISJ}}^{\prime}}_{n}^{k}\right)=\binom{n}{k}$. Since $2{\operatorname{\mathsf{DISJ}}^{\prime}}_{n}^{k}=J-\operatorname{\mathsf{DISJ}}$, where $J$ is the all-ones matrix, we have $\operatorname{\mathrm{rank}}\left({\operatorname{\mathsf{DISJ}}^{\prime}}_{n}^{k}\right)\leq\operatorname{\mathrm{rank}}\left(J\right)+\operatorname{\mathrm{rank}}\left(\operatorname{\mathsf{DISJ}}\right)$. Thus, $\operatorname{\mathrm{rank}}\left(\operatorname{\mathsf{DISJ}}\right)\geq\operatorname{\mathrm{rank}}\left({\operatorname{\mathsf{DISJ}}^{\prime}}_{n}^{k}\right)-\operatorname{\mathrm{rank}}\left(J\right)=\binom{n}{k}-1$. Since $\operatorname{\mathsf{EQ}}$ is an identity matrix, we have $\operatorname{\mathrm{rank}}\left({\operatorname{\mathsf{EQ}}^{\prime}}_{n}^{k}\right)=\binom{n}{k}$. Similar to $\operatorname{\mathsf{DISJ}}$, we have $\operatorname{\mathrm{rank}}\left(\operatorname{\mathsf{EQ}}\right)\geq\operatorname{\mathrm{rank}}\left({\operatorname{\mathsf{EQ}}^{\prime}}_{n}^{k}\right)-1=\binom{n}{k}-1$. ∎
A.3 Approximate Rank
While some properties of approximate rank have been known for real matrices [30, 40], we generalize them to partial matrices. The proofs are adapted from the original proofs.
A.3.1 Error Reduction
Fact A.5 ([30], Fact 1).
Let $10<E<$ 1 be given. Then for each interger $c\geq 1$, there exists a degree-$c$ real univariate polynomial $p(t)$ such that for any $1-E\leq|t|\leq 1+E$,
$|p(t)-\mathrm{sign}(t)|\leq 8\sqrt{c}\left(1-\frac{(1-E)^{2}}{16}\right)^{c},$
where
$\mathrm{sign}(t)\coloneqq\begin{cases}1,&\text{ if }t>0,\\ -1,&\text{ if }t<0.\end{cases}$
Fact A.6 ([33], Lemma 11).
Let $A$ be a real matrix, $p$ be a degree-$d$ polynomial and $B\coloneqq[p(A_{i,j})]_{i,j}$. Then $\operatorname{\mathrm{rank}}\left(B\right)\leq\left(d+1\right)\operatorname{\mathrm{rank}}\left(A\right)^{d}$.
Then we give A.7 and prove it as follow:
Fact A.7 (A generalized version of Corollary 1 in [30]).
Let $F$ be a matrix with $\left\{-1,1,*\right\}$ entries. Let $\epsilon,E$ be constants with $10<\epsilon<E<$ 1. Then $\log\operatorname{\widetilde{rank}}_{\epsilon}(F)=O\left(\log\operatorname{\widetilde{rank}}_{E}(F)\right)$.
Proof.
Let $c$ be any constant positive integer such that
$28\sqrt{c}\left(1-\frac{(1-E)^{2}}{16}\right)^{c}\leq\epsilon/$ 2 .
By A.5, there is a degree-$d$ polynomial $p(t)$ such that for any $1-E\leq|t|\leq 1$,
$|p(t)-\mathrm{sign}(t)|\leq\epsilon/2.$
Let $q(t)=p(t)-\epsilon/2$. Then for any $1-E\leq|t|\leq 1$, we have
$|q(t)-\mathrm{sign}(t)|\leq\epsilon,|q(t)|\leq 1.$
Let $A$ be a real matrix such that $A\in\mathcal{F}_{E}$ and $\operatorname{\mathrm{rank}}\left(A\right)=\operatorname{\widetilde{rank}}_{E}\left(F\right)$, where $\mathcal{F}_{E}$ be the set of real matrices that approximates $F$ with error $E$ as Definition 2.4. Then the matrix $B=[q\left(A_{i,j}\right)]_{i,j}$ satisfies that $B\in\mathcal{F}_{\epsilon}.$ By A.6, we have $\operatorname{\mathrm{rank}}(B)\leq(c+1)\operatorname{\mathrm{rank}}(A)^{c}$. Thus,
$\operatorname{\widetilde{rank}}_{\epsilon}(F)\leq\operatorname{\mathrm{rank}}(B)\leq(c+1)\operatorname{\mathrm{rank}}(A)^{c}=(c+1)\operatorname{\widetilde{rank}}_{E}(F)^{c}.$
Thus, we have $\log\operatorname{\widetilde{rank}}_{\epsilon}(F)=O\left(\log\operatorname{\widetilde{rank}}_{E}(F)\right)$. ∎
A.3.2 Lower Bound via Pattern Matrix Method
We first give some useful definitions, facts and lemmas.
Definition A.8 (Norm of matrices).
For a matrix $A\in\mathbb{R}^{m\times n}$, let the singular values of $A$ be $\sigma_{1}(A)\geq\sigma_{2}(A)\geq\cdots\geq\sigma_{\min\left(m,n\right)}(A)\geq 0$. The spectral norm and trace norm are given by
$||A||$ $\coloneqq\sigma_{1}(A),$
$||A||_{\Sigma}$ $\coloneqq\sum_{i}\sigma_{i}(A).$
Definition A.9 (Approximate trace norm).
For any matrix $F\in\left\{-1,1,*\right\}^{m\times n}$ and $10\leq\epsilon<$ 1, the $\epsilon$-approximate trace norm of $F$ is
$||F||_{\Sigma,\epsilon}\coloneqq\min_{A\in\mathcal{F}_{\epsilon}}||A||_{\Sigma},$
where $\mathcal{F}_{\epsilon}$ is defined in Definition 2.4.
Definition A.10.
For any incomplete matrix $F\in\left\{-1,1,*\right\}^{m\times n}$ and $10\leq\epsilon<$ 1, we say a real matrix $A$ is a real representation of $F$ if:
  1. (1)
    $A_{i,j}=F_{i,j}$ for any $i\in[m],j\in[n]$ such that $F_{i,j}\neq*$;
  2. (2)
    $|A_{i,j}|\leq 1$ for all $i\in[m],j\in[n]$.
Let $\mathcal{F}$ be the set of all real representations of $F$.
By Definitions A.10 and A.9, we have $||F||_{\Sigma,\epsilon}=\min_{A\in\mathcal{F}}||A||_{\Sigma,\epsilon}$.
Fact A.11 ([40], Proposition 2.2).
Let $F\in\mathbb{R}^{m\times n}$ and $\epsilon\geq 0$. Then
$||F||_{\Sigma,\epsilon}\geq\sup_{\Psi\in\mathbb{R}^{m\times n},\Psi\neq 0}\frac{\langle F,\Psi\rangle-\epsilon||\Psi||_{1}}{||\Psi||}.$
As a corollary of A.11, we have the following lemma:
Lemma A.12.
Let $F\in\left\{-1,1,*\right\}^{m\times n}$ and $10\leq\epsilon<$ 1. Then
$||F||_{\Sigma,\epsilon}\geq\sup_{\Psi\in\mathbb{R}^{m\times n},\Psi\neq 0}\frac{\sum_{(x,y)\in\mathrm{dom}}\Psi_{x,y}F_{x,y}-\sum_{(x,y)\notin\mathrm{dom}}|\Psi_{x,y}|-\epsilon||\Psi||_{1}}{||\Psi||}.$
Proof.
By A.11, we have
$||F||_{\Sigma,\epsilon}$ $=\min_{A\in\mathcal{F}}||A||_{\Sigma,\epsilon}$
$\geq\min_{A\in\mathcal{F}}\sup_{\Psi\in\mathbb{R}^{m\times n},\Psi\neq 0}\frac{\langle A,\Psi\rangle-\epsilon||\Psi||_{1}}{||\Psi||}$
$=\sup_{\Psi\in\mathbb{R}^{m\times n},\Psi\neq 0}\frac{\sum_{(x,y)\in\mathrm{dom}}\Psi_{x,y}F_{x,y}-\sum_{(x,y)\notin\mathrm{dom}}|\Psi_{x,y}|-\epsilon||\Psi||_{1}}{||\Psi||}.$
Lemma A.13.
Let $F\in\left\{-1,1,*\right\}^{m\times n}$ and $10\leq\epsilon<$ 1. Then
$\operatorname{\widetilde{rank}}_{\epsilon}(F)\geq\frac{||F||_{\Sigma,\epsilon}^{2}}{\left(1+\epsilon\right)^{2}mn}.$
Proof.
Same as Proposition 2.3 in [40], for any real matrix $A$ that approximates $F$ with $\epsilon$, we have
$||F||_{\Sigma,\epsilon}=\min_{A\in\mathcal{F}_{\epsilon}}||A||_{\Sigma}\leq||A||_{F}\sqrt{\operatorname{\mathrm{rank}}{A}}\leq\sqrt{\sum_{i,j}A_{i,j}^{2}}\sqrt{\operatorname{\mathrm{rank}}{A}}\leq\sqrt{\left(1+\epsilon\right)^{2}mn}\sqrt{\operatorname{\mathrm{rank}}{A}}.$
Thus,
$\operatorname{\widetilde{rank}}_{\epsilon}(F)\geq\operatorname{\mathrm{rank}}{A}\geq\frac{||F||_{\Sigma,\epsilon}^{2}}{\left(1+\epsilon\right)^{2}mn}.$
Fact A.14 ([9], Theorem 10).
Let $F$ be the $(n,t,f)$-pattern matrix, where $f:\{0,1\}^{t}\rightarrow\left\{-1,1,*\right\}$ is given. Suppose $d=\operatorname{\mathsf{\widetilde{deg}}}_{\epsilon}(f)$. Then there exists $\Psi$ be a $(n,t,2^{-n}(n/t)^{-t}\psi)$-pattern matrix for some function $\psi:\{0,1\}^{t}\rightarrow\mathbb{R}$ such that
$||\Psi||_{1}$ $=1,$ (5)
$\sum_{(x,y)\in\mathrm{dom}}F_{x,y}\Psi_{x,y}-\sum_{(x,y)\notin\mathrm{dom}}|\Psi_{x,y}|$ $>\epsilon,$
$||\Psi||$ $\leq\left(\frac{t}{n}\right)^{d/2}\left(2^{n+t}\left(\frac{n}{t}\right)^{t}\right)^{-1/2}.$
Finally, we give the following lower bound of approximate rank via pattern matrix method by A.15.
Fact A.15 (A generalized version of Theorem 1.4 in [40]).
Let $F$ be the $(n,t,f)$-pattern matrix, where $f:\{0,1\}^{t}\rightarrow\left\{-1,1,*\right\}$ is given. Then for every $\epsilon\in[0,1)$ and every $\delta\in[0,\epsilon]$,
$\operatorname{\widetilde{rank}}_{\delta}(F)\geq\left(\frac{\epsilon-\delta}{1+\delta}\right)^{2}\left(\frac{n}{t}\right)^{\operatorname{\mathsf{\widetilde{deg}}}_{\epsilon}(f)}.$
Proof.
By Lemma A.12 and Equation 5, we have
$||F||_{\Sigma,\delta}\geq(\epsilon-\delta)\left(\frac{n}{t}\right)^{\operatorname{\mathsf{\widetilde{deg}}}_{\epsilon}(f)/2}\sqrt{2^{n+t}(n/t)^{t}}.$ (6)
Since $F$ is a $2^{n}\times(n/t)^{t}2^{t}$ matrix, by Lemma A.13 and Equation 6, we have
$\operatorname{\widetilde{rank}}_{\delta}(F)$ $\geq\frac{||F||_{\Sigma,\delta}^{2}}{\left(1+\delta\right)^{2}2^{n+t}(n/t)^{t}}$
$\geq\left(\frac{\epsilon-\delta}{1+\delta}\right)^{2}\left(\frac{n}{t}\right)^{\operatorname{\mathsf{\widetilde{deg}}}_{\epsilon}(f)}.$
A.4 Sampling in Communication Model
First, we give Fact A.16. Then we restate and prove its randomized version, Fact 3.12.
Fact A.16 ([15], Proposition 2).
Suppose $x,y\in\{0,1\}^{n}$ are inputs of Alice and Bob such that $|x|\neq|y|$. Alice and Bob can find an index $i$ such that $x_{i}\neq y_{i}$ using $O(\log n)$ bits of communication.
See 3.12
Proof.
The protocol of A.16 is a deterministic protocol based on binary search. Initially, the search space is $[n]$. In each round, Alice and Bob shrink the search space to one-half of the original search space. Finally, Alice and Bob find one index $i\in[n]$ such $x_{i}\neq y_{i}$ by $O(\log n)$ rounds. The deterministic protocol can be adjusted to a randomized protocol easily. By making the same random permutation to the bits of the inputs using public coins, Alice and Bob can find an index $i$ such that $x_{i}\neq y_{i}$ uniformly. ∎
A.5 Quantum Amplitude Amplification
Fact A.17 ([27], Merging Theorem 3 and item (4) in Corollary 4).
Given a unitary $U$ such that $U\ket{0}=\ket{\psi}$ and an orthogonal projector $\Pi$, there exists a quantum algorithm outputting an estimation $\tilde{p}$ of $p=\langle\psi|\Pi|\psi\rangle$ such that
$|\tilde{p}-p|\leq\frac{1}{3}\epsilon$
by $O\left(\sqrt{p}/\epsilon\right)$ calls to (the controlled versions of) $U,U^{{\dagger}}$ and $I-2\Pi$ with the bounded error.
As a direct corollary of A.17, we give the quantum version of 3.11 as follows.
Fact A.18.
Given a set $S$, suppose there is a subset $S^{\prime}$ of $S$ such that $\frac{|S^{\prime}|}{|S|}\leq\beta-\epsilon$ or $\frac{|S^{\prime}|}{|S|}\geq\beta+\epsilon$. Suppose we have a quantum sampler such that $\mathcal{S}\ket{0}=\frac{1}{\sqrt{|S|}}\sum_{i\in S}\ket{i}$. Let orthogonal projector $\Pi$ be defined as $\Pi=\sum_{i\in S^{\prime}}\ket{i}\bra{i}$ |. Since $\frac{|S^{\prime}|}{|S|}=\bra{0}\mathcal{S}^{{\dagger}}\Pi\mathcal{S}\ket{0}$, we can decide whether $\frac{|S^{\prime}|}{|S|}\leq\beta-\epsilon$ or $\frac{|S^{\prime}|}{|S|}\geq\beta+\epsilon$ by $O\left(\sqrt{\beta}/\epsilon\right)$ calls of (the controlled versions of) $\mathcal{S},\mathcal{S}^{{\dagger}}$ and $I-2\Pi$ with the bounded error.
Remark A.19.
[…]
Appendix B Quantum and Randomized Communication Complexity of Permutation-Invariant Functions
B.1 The Lower Bound on Quantum Communication Complexity
In this section, we restate and prove the following lemmas: See 3.7
Proof.
  1. 1.
    Suppose there exists a quantum protocol $\mathcal{P}$ that solves $\mathsf{ESetInc}_{a+\ell_{1}+\ell_{3},b+\ell_{2}+\ell_{3},c+\ell_{3},g}^{n+\ell}$. To solve $\mathsf{ESetInc}_{a,b,c,g}^{n}$, Alice and Bob can append some bit strings to their input, respectively. First, Alice appends $\ell_{1}$ 1’s to her input string, and Bob appends $\ell_{1}$ 0’s in the same positions. Second, Bob appends $\ell_{2}$ 1’s to his input, and Alice appends $\ell_{2}$ 0’s in the same positions. Third, Alice and Bob append $\ell_{3}$ 1’s at the same time. Finally, Alice and Bob run the protocol for $\mathsf{ESetInc}_{a+\ell_{1}+\ell_{3},b+\ell_{2}+\ell_{3},c+\ell_{3},g}^{n+\ell}$.
  2. 2.
    Suppose there exists a quantum protocol solving $\mathsf{ESetInc}_{a,b,c,g}^{n}$. To solve $\mathsf{ESetInc}_{a,n-b,a-c,g}^{n}$, Bob can first flip each bit of his input, and then Alice and Bob run the protocol for $\mathsf{ESetInc}_{a,b,c,g}^{n}$. Thus $Q^{*}\left(\mathsf{ESetInc}_{a,n-b,a-c,g}^{n}\right)\leq Q^{*}\left(\mathsf{ESetInc}_{a,b,c,g}^{n}\right)$. By a similar argument, we have
    $Q^{*}\left(\mathsf{ESetInc}_{a,b,c,g}^{n}\right)$ $\leq Q^{*}\left(\mathsf{ESetInc}_{a,n-b,a-c,g}^{n}\right),$
    $Q^{*}\left(\mathsf{ESetInc}_{a,b,c,g}^{n}\right)$ $\leq Q^{*}\left(\mathsf{ESetInc}_{n-a,b,b-c,g}^{n}\right),$
    $Q^{*}\left(\mathsf{ESetInc}_{n-a,b,b-c,g}^{n}\right)$ $\leq Q^{*}\left(\mathsf{ESetInc}_{a,b,c,g}^{n}\right).$
    Thus, $Q^{*}\left(\mathsf{ESetInc}_{a,b,c,g}^{n}\right)=Q^{*}\left(\mathsf{ESetInc}_{a,n-b,a-c,g}^{n}\right)=Q^{*}\left(\mathsf{ESetInc}_{n-a,b,b-c,g}^{n}\right)$.
  3. 3.
    Suppose there exists a quantum protocol solving $\mathsf{ESetInc}_{ka,kb,kc,kg}^{kn}$. To solve $\mathsf{ESetInc}_{a,b,c,g}^{n}$, Alice and Bob can repeat their input strings $k$ times and run the protocol for $\mathsf{ESetInc}_{ka,kb,kc,kg}^{kn}$.
See 3.8
Proof.
We consider the function $f_{k,l}:\{0,1\}^{k}\rightarrow\left\{-1,1,*\right\}$ given by
$f_{k,l}(x)\coloneqq\begin{cases}-1,&\text{if $|x|=l-1/2,$}\\ 1,&\text{if $|x|=l+1/2,$}\\ *,&\text{otherwise.}\end{cases}$ (7)
Let $D$ be a Boolean predicate such that $D(|x|)=f(x)$ for any $x\in\{0,1\}^{k}$. By A.3, we have
$\Gamma(D)=|2(l-1/2)-k+1|=|2l-k|=k-2l,$
and thus
$\operatorname{\mathsf{\widetilde{deg}}}_{1/3}\left(f_{k,l}\right)=\Omega\left(\sqrt{k(k-\Gamma(D))}\right)=\Omega\left(\sqrt{kl}\right).$ (8)
Let $P$ be the $\left(2k,k,f_{k,l}\right)$-pattern matrix defined as Definition A.1. A.4 implies that $Q^{*}_{1/3}\left(P\right)=\Omega\left(\sqrt{kl}\right)$, where
$P$ $=\Big{[}f_{k,l}(x|_{V}\oplus w)\Big{]}_{x\in\{0,1\}^{2k},\,(V,w)\in\mathcal{V}(2k,k)\times\{0,1\}^{k}}$
$=\Big{[}f_{k,l}(x_{1}\overline{x}_{1}x_{2}\overline{x}_{2}\ldots x_{2k}\overline{x_{2k}}|_{V})\Big{]}_{x\in\{0,1\}^{2k},\,V\in\mathcal{V}(4k,k)}.$
For any $x\in\{0,1\}^{2k}$, we have $x_{1}\overline{x}_{1}x_{2}\overline{x}_{2}\ldots x_{2k}\overline{x_{2k}}\in\{0,1\}^{4k}$ and $|x_{1}\overline{x}_{1}x_{2}\overline{x}_{2}\ldots x_{2k}\overline{x_{2k}}|=2k$; for any $V\in\mathcal{V}(4k,k)$, we have $V\in\{0,1\}^{4k}$ and $|V|=k$. Thus, $P$ is a submatrix of $\mathsf{ESetInc}^{4k}_{2k,k,l,1/2}$, which is defined as
$\mathsf{ESetInc}^{4k}_{2k,k,l,1/2}(x,y)\coloneqq\begin{cases}-1,&\text{if $|x|=2k,|y|=k$ and $|x\land y|=l-1/2$},\\ 1,&\text{if $|x|=2k,|y|=k$ and $|x\land y|=l+1/2$},\\ *,&\text{otherwise.}\end{cases}$
As a result, we have $Q^{*}_{1/3}\left(\mathsf{ESetInc}^{4k}_{2k,k,l,1/2}\right)\geq Q^{*}_{1/3}\left(P\right)=\Omega\left(\sqrt{kl}\right)$. ∎
B.2 Quantum Upper Bound of Set-Inclusion Problem
We restate and prove Lemma 3.10 as follows: See 3.10
Proof.
The proof of Lemma 3.10 is similar to the proof of Lemma 3.9. The only essential difference is that 3.11 is replaced by its quantum speedup version, A.18. For simplicity, we discuss the following two cases as the proof of Lemma 3.9. For other cases, we can obtain the same result similarly. In the following proof, for $x\in\{0,1\}^{n}$, let $S_{x}\coloneqq\left\{i:x_{i}=1\right\}$. Let $\Pi\coloneqq\sum_{i\in S_{x\land y}}\ket{i}\bra{i}$ |.
  1. 1.
    $n_{1}=c,n_{2}=a-c$. In this case, the problem to compute $\mathsf{SetInc}_{a,b,c,g}^{n}(x,y)$ can be reduced equivalently to estimate $\frac{|x\land y|}{|x|}$. We first prove some special unitary operations can be constructed with $O(\log n)$ qubits communication.
    1. (a)
      Since Alice knows all information about $x$, she can implement $S_{x}$ such that $\mathcal{S}_{x}\ket{0}=\frac{1}{\sqrt{a}}\sum_{i\in S_{x}}\ket{i}$ by herself without communication. Let |$\ket{\phi_{1}}\coloneqq\mathcal{S}_{x}\ket{0}$. Then we have
      $\bra{\phi_{1}}\Pi\ket{\phi_{1}}=\frac{|x\land y|}{|x|}.$
    2. (b)
      For inputs $x,y\in\{0,1\}^{n}$ and any $i\in[n]$, let unitary operator $O_{x},O_{y},O_{x\land y}$ be defined as
      $O_{x}\ket{i}$ $\coloneqq(-1)^{x_{i}}\ket{i},$
      $O_{y}\ket{i}$ $\coloneqq(-1)^{y_{i}}\ket{i},$
      $O_{x\land y}\ket{i}$ $\coloneqq(-1)^{x_{i}y_{i}}\ket{i}.$
      Then $I-2\Pi=O_{x\land y}$. For any |$\ket{\phi}\coloneqq\sum_{i\in[n]}\alpha_{i}\ket{i}$, Alice and Bob can perform $I-2\Pi$ to any state |$\ket{\phi}$ using $O(\log n)$ qubits of quantum communication as follows: i) Alice performs $O_{x}$ to |$\ket{\phi}$ and sends |$\ket{\phi^{\prime}}=O_{x}\ket{\phi}=\sum_{i\in[n]}\alpha_{i}(-1)^{x_{i}}\ket{i}$ to Bob; ii) Bob performs $O_{y}$ to |$\ket{\phi^{\prime}}$ and obtains $\sum_{i\in[n]}\alpha_{i}(-1)^{x_{i}y_{i}}\ket{i}=O_{x\land y}\ket{\phi}$.
    By A.18, a quantum speedup version of 3.11, and using an argument similar to Lemma 3.9, we can obtain the following conclusion: using $O\left(\frac{\sqrt{n_{1}n_{2}}}{g}\right)$ unitary operations $\mathcal{S}_{x}$ and $I-2\Pi$, Alice and Bob can estimate $\frac{|x\land y|}{|x|}$ with errors at most $O\left(\frac{g}{a}\right)$. Furthermore, they can decide whether $|x\land y|\geq c+g$ or $|x\land y|\leq c-g$ with success probability at least $1-1/(6\log n)$ using $O\left(\frac{\sqrt{n_{1}n_{2}}}{g}\log n\log\log n\right)$ qubits communication.
  2. 2.
    $n_{1}=c$, $n_{2}=n-a-b+c$ and $a+b<n$. In this case, the problem to compute $\mathsf{SetInc}_{a,b,c,g}^{n}(x,y)$ can be reduced equivalently to estimate $\frac{|x\land y|}{|\overline{x}\oplus y|}$. Similar to Case 1, we first prove Alice and Bob can construct a quantum sampler $\mathcal{S}_{\overline{x}\oplus y}$ such that
    $\mathcal{S}_{\overline{x}\oplus y}\ket{0}\ket{0}=\frac{1}{\sqrt{|N|}}\sum_{f\in[N]}\ket{ff}\ket{f(x,y)}=\ket{\phi_{2}},$
    using $O\left(\log n\right)$ qubits communication, where $N=C\cdot\frac{n_{1}n_{2}}{g^{2}}$ for some enough large constant $C$ and each $f$ encodes a random function such that $f(x,y)$ is a uniform distribution of the elements in $S_{\overline{x}\oplus y}$. Specifically, the construction procedure is as follows: First, Alice and Bob use $O\left(\log\frac{n_{1}n_{2}}{g^{2}}\right)$ qubits communication to transform |$\ket{0}$ into a maximum entanglement bipartite state $\frac{1}{\sqrt{N}}\sum_{f\in[N]}\ket{ff}$. Moreover, both Alice and Bob hold one part of the state. Second, by Fact 3.12, Alice and Bob can sample an element from $S_{\overline{x}\oplus y}$ uniformly using $O(\log n)$ bits of communication. Equivalently, Alice and Bob can generate a random function $f$ using public coins, and output $f(x,y)$ using $O(\log n)$ bits of communication such that the distribution of the output is a uniform distribution of the elements in $S_{\overline{x}\oplus y}$. Since quantum circuits can simulate classical circuits efficiently, Alice and Bob can perform a uniform operation that transforms |$\ket{ff}\ket{b}$ into |$\ket{ff}\ket{f(x,y)+b}$ for any $f\in N$ using $O\left(\log n\right)$ qubits communication, where both Alice and Bob hold one part of |$\ket{ff}$, $b\in\left\{0,1,...,n-1\right\}$ and the addition is with modulo $n$. As a whole, Alice and Bob can construct the above quantum sampler $\mathcal{S}_{\overline{x}\oplus y}$ using $O\left(\log\frac{n_{1}n_{2}}{g^{2}}+\log n\right)=O\left(\log n\right)$ qubits communication. Furthermore, since $f(x,y)$ is a uniform distribution of the elements in $S_{\overline{x}\oplus y}$ for any $f\in\left\{0,1,...,N-1\right\}$, $N=C\cdot\frac{n_{1}n_{2}}{g^{2}}$ for some enough large constant $C$, and
    $\bra{\phi_{2}}\left(I\otimes\Pi\right)\ket{\phi_{2}}=\frac{|\left\{f:f(x,y)\in S_{x\land y}\right\}|}{N},$
    we have
    $|\bra{\phi_{2}}\left(I\otimes\Pi\right)\ket{\phi_{2}}-\frac{|x\land y|}{|\overline{x}\oplus y|}|=O\left(\frac{g}{m}\right)$ (9)
    similar to the analysis of Case 1 of Lemma 3.9, where $m=n_{1}+n_{2}$.
By A.18, a quantum speedup version of 3.11, and using an argument similar to Lemma 3.9, we can obtain the following conclusion, using $O\left(\frac{\sqrt{n_{1}n_{2}}}{g}\right)$ unitary operations $\mathcal{S}_{|\overline{x}\oplus y|}$ and $I-2\Pi$, Alice and Bob can estimate $\bra{\phi_{2}}\left(I\otimes\Pi\right)\ket{\phi_{2}}$ with errors at most $O\left(\frac{g}{m}\right)$. By Equation (9), Alice and Bob also can estimate $\frac{|x\land y|}{|\overline{x}\oplus y|}$ with errors at most $O\left(\frac{g}{m}\right)$. Next, using the same argument as Lemma 3.9, we can prove that Alice and Bob can decide whether $|x\land y|\geq c+g$ or $|x\land y|\leq c-g$ with success probability at least $1-1/(6\log n)$ using $O\left(\frac{\sqrt{n_{1}n_{2}}}{g}\log n\log\log n\right)$ qubits communication. ∎
Appendix C Log-Rank Conjecture of Permutation-Invariant Functions
In this section, we restate and prove Lemma 4.2. See 4.2
Proof.
We give the following deterministic protocol. First, Alice and Bob exchange the values of $|x|,|y|$ by $O(\log n)$ bits. Suppose $|x|=a,|y|=b$. i) If $f_{a,b}=c$ for some constant number $c\in\left\{-1,1\right\}$, then they output $c$ directly. ii) If $f_{a,b}$ is not a constant function, then they perform the following operations: if $\binom{n}{a}\leq\binom{n}{b}$, Alice sends $x$ to Bob using $\log\binom{n}{a}$ bits, and then Bob outputs $f_{a,b}(x,y)$; otherwise, Bob sends $y$ to Alice using $\log\binom{n}{b}$ bits, and then Alice outputs $f_{a,b}(x,y)$. In total, the communication cost of the protocol is
$\log n+\max_{a,b:f_{a,b}\text{ is not constant}}\min\left\{\log\binom{n}{a},\log\binom{n}{b}\right\}$ (10)
on the worst case. If $f_{a,b}$ is not a constant function, then $0<a,b<n$, and thus
$\log n=\log\binom{n}{1}\leq\min\left\{\log\binom{n}{a},\log\binom{n}{b}\right\}.$ (11)
If $a\leq n/2$, then $\log\binom{n}{a}\leq a\log n$; if $a>n/2$, then $\log\binom{n}{a}=\log\binom{n}{n-a}\leq(n-a)\log n$. Thus,
$\log\binom{n}{a}\leq\min\left\{a,n-a\right\}\cdot\log n.$ (12)
Combining Equations 10, 11 and 12, the communication cost of the protocol is
$O\left(\max_{a,b:f_{a,b}\text{ is not constant}}\min\left\{a,b,n-a,n-b\right\}\cdot\log n\right).$
Appendix D Log-Approximate-Rank Conjecture of Permutation-Invariant Functions
In this section, we restate and prove Lemmas 5.1 and 5.2. See 5.1
Proof.
By the definition of approximate rank (Definition 2.4), to prove $\operatorname{\widetilde{rank}}(A)\leq\operatorname{\widetilde{rank}}(B)$, it suffices to prove $A$ is a submatrix of $B$. Then we discuss three cases:
  1. 1.
    For $x,y\in\{0,1\}^{n}$, let
    $x^{\prime}$ $\coloneqq x\underbrace{1\cdots 1}_{l_{1}}\underbrace{0\cdots 0}_{l_{2}}\underbrace{1\cdots 1}_{l_{3}}\underbrace{0\cdots 0}_{l-\left(l_{1}+l_{2}+l_{3}\right)},$
    $y^{\prime}$ $\coloneqq y\underbrace{0\cdots 0}_{l_{1}}\underbrace{1\cdots 1}_{l_{2}}\underbrace{1\cdots 1}_{l_{3}}\underbrace{0\cdots 0}_{l-\left(l_{1}+l_{2}+l_{3}\right)}.$
    Then
    $\mathsf{ESetInc}_{a,b,c,g}^{n}\left(x,y\right)=\mathsf{ESetInc}_{a+\ell_{1}+\ell_{3},b+\ell_{2}+\ell_{3},c+\ell_{3},g}^{n+\ell}\left(x^{\prime},y^{\prime}\right),$
    Thus, $\mathsf{ESetInc}_{a,b,c,g}^{n}$ is a submatrix of $\mathsf{ESetInc}_{a+\ell_{1}+\ell_{3},b+\ell_{2}+\ell_{3},c+\ell_{3},g}^{n+\ell}$.
  2. 2.
    For $x,y\in\{0,1\}^{n}$, we have
    $\mathsf{ESetInc}_{a,b,c,g}^{n}(x,y)=\mathsf{ESetInc}_{a,n-b,a-c,g}^{n}(x,\overline{y})=\mathsf{ESetInc}_{n-a,b,b-c,g}^{n}(\overline{x},y).$
  3. 3.
    Since $\mathsf{ESetInc}_{a,b,c,g}^{n}(x,y)=\mathsf{ESetInc}_{ka,kb,kc,kg}^{kn}(\underbrace{x\cdots x}_{k},\underbrace{y\cdots y}_{k})$, we have $\mathsf{ESetInc}_{a,b,c,g}^{n}$ is a submatrix of $\mathsf{ESetInc}_{ka,kb,kc,kg}^{kn}$.
See 5.2
Proof.
Let $P$ be the $\left(2k,k,f_{k,l}\right)$-pattern matrix, where $f_{k,l}$ is defined in Equation 7. By Lemma 3.8, $P$ is a submatrix of $\mathsf{ESetInc}^{4k}_{2k,k,l,1/2}$. Thus, we have
$\log\left(\operatorname{\widetilde{rank}}\left(\mathsf{ESetInc}^{4k}_{2k,k,l,1/2}\right)\right)\geq\log\left(\operatorname{\widetilde{rank}}\left(P\right)\right).$
By Facts A.7 and A.15 and Equation 8, we have
$\log\left(\operatorname{\widetilde{rank}}\left(P\right)\right)=\Omega\left(\operatorname{\mathsf{\widetilde{deg}}}\left(f_{k,l}\right)\right).$
Thus,
$\log\left(\operatorname{\widetilde{rank}}\left(\mathsf{ESetInc}^{4k}_{2k,k,l,1/2}\right)\right)=\Omega\left(\operatorname{\mathsf{\widetilde{deg}}}\left(f_{k,l}\right)\right)=\Omega\left(\sqrt{kl}\right),$
where the second equality comes from Equation 8. ∎